### v1

# Fast Dynamic 3D Object Generation from a Single-view Video

Abstract
Generating dynamic three-dimensional (3D) object from a single-view video is challenging due to the lack of 4D labeled data. Existing methods extend text-to-3D pipelines by transferring off-the-shelf image generation models such as score distillation sampling, but they are slow and expensive to scale (e.g., 150 minutes per object) due to the need for back-propagating the information-limited supervision signals through a large pretrained model. To address this limitation, we propose an efficient video-to-4D object generation framework called Efficient4D. It generates high-quality spacetime-consistent images under different camera views, and then uses them as labeled data to directly train a novel 4D Gaussian splatting model with explicit point cloud geometry, enabling real-time rendering under continuous camera trajectories. Extensive experiments on synthetic and real videos show that Efficient4D offers a remarkable 10-fold increase in speed when compared to prior art alternatives while preserving the same level of innovative view synthesis quality. For example, Efficient4D takes only 14 minutes to model a dynamic object.
11footnotetext:  Corresponding to lizhangfd@fudan.edu.cn
1 Introduction
Humans possess a remarkable capacity to comprehensively comprehend the spatial and temporal characteristics of a dynamic object in a brief video, even with a limited perspective, enabling them to predict its appearance in unseen viewpoints over time. Despite the significant advancement of three-dimensional (3D) object generation, existing works [28, 17, 4, 41] mostly consider static scenes or objects. With the availability of large-scale 3D datasets [6, 5], training generalizable models capable of directly generating multi-view images becomes possible [20, 21, 22]. These generated images can be turned into a 3D object through reconstruction techniques [25, 39]. By further augmenting these generated static objects with predefined animations [1], dynamic 3D content can be generated. However, this is highly limited due to the need for fine-grained meshes as well as rigid restrictions.
Refer to caption Refer to caption
Refer to caption Refer to caption
Input Generated dynamic objects
Figure 1: Examples of video-to-4D generation. Input: A brief video of a dynamic object, as represented by 3 frames per case; Output: Generated novel views at different timestamps.
Directly generating 4D object/scene content from text description has been recently attempted [36]. To bypass the need of exhaustively labeled training data pairs in form of (text, 4D), it trains a Neural Radiance Fields (NeRF)-like representation [25] via score distillation sampling [28] and separates the modeling of static scene and its dynamics. Not only is this method computationally inefficient caused by heavy supervision back-propagation through a large pretrained model, but also its textual condition is highly ambiguous in expressing the intended visual content. In quest of the aforementioned human’s capability, a recent work [10] proposed to generate dynamic 3D object images from a single-view video (statically captured monocular video from a fixed view), namely as video-to-4D object generation. However, similar as [36] this method is also slow to train (e.g., 150 minutes to model a single dynamic object) in addition to complex design, hence unscalable in practice.
To address identified limitations, we formulate an efficient video-to-4D object generation method in a two-staged pipeline called Efficient4D. In the first stage, we generate spacetime-consistent images across different camera views as synthetic training data. This is realized by imposing temporal smoothing filter into a previous multi-view image generator (e.g., SyncDreamer [21]) in tandem with frame interpolation. In the second stage, we use these training data to optimize a novel 4D Gaussian splatting model. This is an extension of the recent 3D Gaussian splatting [12], originally designed for static 3D scene representation, with the temporal dimension introduced additionally on top. Using Gaussian representation brings about further computational efficiency gain along with the explicit geometry property, when compared with NeRF based designs. To tackle the discontinuity between generated frames, we design a inconsistency-aware loss function. A couple of video-to-4D results are presented in Fig. 1.
Our contributions are summarized as follows: (i) We consider for the first time the efficiency challenge with the under-studied video-to-4D object generation problem. (ii) We propose an efficient video-to-4D object generation pipeline, Efficient4D, characterized by directly generating high-quality training data without the need for heavy supervision back-propagation through a large pretrained model as suffered by most 3D/4D object generation approaches. (iii) Instead of using NeRF as the 3D representation model, we introduce a novel 4D Guardian splatting model with the real-time rendering advantage.
(iv) Extensive experiments on both synthetic and real videos validate the significant efficiency advantage (e.g., 10$\times$ speedup) of our Efficient4D over the prior art, whilst maintaining the same level of new view synthesis. Additionally, our method can also work well under the more challenging few-shot setting where only a handful of key frames are available for training, further extending the application scope.
2 Related work
4D representation
Efforts to synthesize videos with free viewpoint control in dynamic scenes have a well-documented history [46]. For example, pre-NeRF [25] approaches faced challenges in reconstructing intricate scene details. Recent advancements in 4D representations, particularly those based on neural rendering, include D-NeRF [29], which decouples geometry and motion, utilizing a canonical space and a learned deformation field. HyperNeRF [27] extends the deformation field to a higher dimension, employing the level-set method for smoother optimization. DeVRF [18] reconstructs the canonical space from multi-view dense observations in the initial frame and learns the deformation field with sparse views in subsequent timesteps, utilizing a voxel grid for representation.
Similarly, deformable 3D Gaussians [44] integrates the deformation field with 3D Gaussian Splatting for the scene’s geometric structure. DyNeRF [14] extends NeRF with an additional latent code to capture detailed temporal variations, while DynIBaR [16] extends the image-based rendering paradigm for representing long videos with complex camera and object motions. Another group of method [7, 3, 33] adopts tensor decomposition of 4D volumes to represent dynamic 3D scenes.
However, these representations often involve complex training schedules or suffer from slow convergence. In this work, drawing inspiration from 3D Gaussian Splatting [12], we propose an efficient 4D scene primitive for the representation of dynamic scenes.
3D generation
3D generation takes two main settigns: text-to-3D and image-to-3D. The pioneering work, DreamFusion [28], introduces the score distillation sampling (SDS) loss for optimizing 3D shapes with diffusion models. SDS’s generality has prompted numerous subsequent efforts in both text-to-3D tasks [17, 24, 32, 34, 4, 15, 38, 8, 42, 41, 45] and image-to-3D tasks [43, 37, 23, 30] across various dimensions.
However, SDS-based approaches often suffer from extended optimization times. Conversely, some efficient methods [26, 11, 20, 19, 21] have emerged. Notably, Point-E [26] and Shap-E [11] train models to directly generate 3D point clouds or meshes. Zero123 [20] focuses on generating a 2D image from an unseen view based on a single image, convertible to a 3D shape through SDS or [19]. Importantly, SyncDreamer [21] produces multi-view consistent images, offering inspiration for reconstructing 3D objects.
4D generation
There are a few recent works dedicated for more challenging 4D object generation. For instance, MAV3D [36] deals with a text-to-4D problem by training Hexplane [3] with a video diffusion model and SDS loss. Instead of text input, Consistent4D [10] conditions the generation of 4D object over time on a monocular video with richer and more specific information. However, this work is computationally inefficient due to inheriting the previous SDS loss, along with complex pipeline design. To overcome this limitation, we present a novel two-staged pipeline in a generation-and-reconstruction strategy, drastically boosting the model training speed by 10$\times$ whilst maintaining the quality of novel view synthesis.
Figure 2: Overview of our Efficient4D approach. Given as the input a brief video depicting a dynamic object in limited perspectives, our model aims to generate this object with geometrical and temporal consistency under any specific view and time. Efficient4D comprises two components: (A) Image sequence synthesis across views and timestamps, resulting in (c) an image matrix where each row consists of multi-view geometrically consistent images and each column consists of view-specific temporally consistent images. (B) A novel 4D Gaussian representation model (d) that represents the scene with a number of Gaussian points. It can be trained efficiently and robustly under the confidence-aware (e) supervision on the generated image matrix.
3 Method
Our Efficient4D addresses the challenge of efficiently generating dynamic objects under novel views from a single-view video . The model, illustrated in Fig. 2, comprises two key components:
  • An image synthesis pipeline (Sec. 3.1) generates images across views and timestamps, ensuring sufficient geometry and temporal consistency.
  • A novel 4D Gaussian representation model (Sec. 3.2) efficiently utilizes the synthetic images for accurate dynamic object reconstruction and novel view synthesis.
3.1 Image synthesis across views and timestamps
Due to the difficulty of obtaining calibrated 4D scans, our approach involves the direct generation of high-quality consistent 4D data from a single-view video which is much easier to capture (Fig. 2(A)). Specifically, we seek to produce a $M\times N$ image matrix $\mathcal{D}=\{e_{ij}\}_{i,j=1}^{M,N}$ representing 2D images with geometrical and temporal consistency. Here, $M$ denotes timestamps, and $N$ represents views, with each matrix element corresponding to an image (Fig. 2(c)). This approach combines conventional video (capturing time variation, represented by a single column in the image matrix) and 3D (capturing view variation, represented by a single row in the image matrix) generation [35, 21], offering comprehensive information for modeling a dynamic object.
To initiate the image matrix $\mathcal{D}$, we set the first view (i.e., the first column) with $K$ frames from the input video (Fig. 2(a)) and proceed to generate the remaining views. Our task involves generating multi-view consistent images from a single image for each row. Existing image-to-3D methods, such as SyncDreamer [21], can be leveraged for this purpose. However, these methods often struggle with temporal inconsistency within a specific view (i.e., continuity in the column direction) due to the independent synthesis of multi-frame images. To address this issue, we propose an enhanced version of SyncDreamer with improved temporal continuity, referred to as SyncDreamer-T.
Specifically, SyncDreamer generates $N$ multi-view images $\{\mathbf{x}_{0}^{(1)},\dots,\mathbf{x}_{0}^{(N)}\}$ of a static object using a synchronized $N$-view noise predictor $\{\epsilon_{\theta}^{(n)}|n=1,\dots,N\}$ that predicts synchronized noise for noisy multi-view images $\mathbf{x}_{t}^{1:N}$. The noise predictor is conditioned on information correlated with all views. Cross-view conditioning is achieved through a spatial feature volume $\mathcal{V}\in\mathbb{R}^{F\times V\times V\times V}$ unprojected by $\mathbf{x}_{t}^{1:N}$ to inject 3D-aware features into the shared noise predictor, ensuring geometrical consistency across views for static moments.
To impose temporal consistency, the information from $\mathcal{V}$ at different timestamps is aggregated using a time-synchronous spatial volume we design here (Fig. 2(b)). A smoothing filter layer is introduced into the spatial volumes of different frames/timestamps, incorporating a weight vector $\mathbf{w}=(w_{-k},\dots,w_{0},\dots,w_{k})$. At each denoising step, time-synchronized spatial volumes $\tilde{\mathcal{V}}_{i}$ for each input frame $i\in\{1,2,\dots,M\}$ are constructed as:
$\tilde{\mathcal{V}}_{i}=\sum_{j=-k}^{k}w_{i+j}\mathcal{V}_{i+j}.$ (1)
This synchronization ensures consistent features across $k$ past and $k$ future frames during the denoising process, enhancing temporal consistency.
To further improve temporal resolution, video frame interpolation (e.g., RIFE [9]) is applied after generating the image matrix $\mathcal{D}$ as described above. The midpoint interpolation is applied twice recursively, giving three additional frames between each pair of consecutive frames. This results in a total of $M=4K-3$ images in a column of $\mathcal{D}$.
Analysis on temporal synchronization
For a clearer insight into our temporal synchronization design, we undertake a simplified experiment involving two input frames, each with its feature volumes labeled as $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$. The fusion process is carried out by combining them as follows:
$\tilde{\mathcal{V}}_{i}=(1-w)\mathcal{V}_{i}+w\mathcal{V}_{3-i},$ (2)
where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion” as illustrated in Fig. 3. For $w=0$, the original generation process is represented, where the spatial volumes of the two frames are independent, leading to temporal inconsistency. As $w$ approaches 0.5, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry. This suggests that smoothing at the feature level is effective in aligning frames over time.
However, a challenge arises as the fused features may induce similar motion. While the bottom-left astronaut is stepping forward, the one at the bottom-right column does not exhibit forward motion like the top astronaut. Thus, a trade-off is necessary between achieving temporal consistency and preserving motion independence. In practical terms, it is recommended to set the ratio between 0.25 and 0.4, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry.
Refer to caption Refer to caption
Input Feature fusion
Figure 3: Example analysis on temporal synchronization. In this illustration, we manipulate the fusion ratio ($w$) across a range from 0 to 0.5 for the spatial feature volumes of two input frames. The results indicate that a moderate ratio can achieve superior outcomes by balancing both temporal consistency and motion independence.
3.2 4D generation through reconstruction
Recall that we desire a 3D dynamic object, we should not be content with the discrete images. Our next goal is to model a truly 4D content from the image matrix $\mathcal{D}$. For efficient modeling, we formulate a novel 4D Gaussian representation model (Fig. 2(B)), departing from previous slow-to-train 4D reconstruction models [7].
4D Gaussian representation model
Our model builds upon the 3D Gaussian splatting technique introduced in [12], originally designed for static scene representation. We extend it to address the complexities of dynamic scenes, enabling the real-time synthesis of high-fidelity views in 4D space. The model represents each Gaussian defined as:
$G(\mathbf{p}|\mathbf{\mu},\Sigma)=e^{-\frac{1}{2}\left(\mathbf{p}-\mathbf{\mu}\right)^{\top}\Sigma^{-1}\left(\mathbf{p}-\mathbf{\mu}\right)},$ (3)
where $\mathbf{\mu}\in\mathbb{R}^{4}$ is the mean vector, and $\Sigma\in\mathbb{R}^{4\times 4}$ is the anisotropic covariance matrix. The input $\mathbf{p}=(\mathbf{x},t)\in\mathbb{R}^{4}$ represents a spacetime position with a spatial coordinate $\mathbf{x}$ and time $t$. The covariance matrix $\Sigma$ decomposes into a diagonal scaling matrix $S\in\mathbb{R}^{4\times 4}$ and a rotation matrix $R\in\mathbb{R}^{4\times 4}$ through $\Sigma=RSS^{\top}R^{\top}$. The 4D rotation $R$ is represented by a pair of iso-rotations, each characterized by a quaternion.
In contrast to the 3D counterpart, our formulation incorporates the time dimension, allowing coherent modeling of dynamics in both space and time.
Rendering
For rendering, each Gaussian includes opacity $\alpha$ and view-dependent color $\mathbf{c}$ represented by spherical harmonics (SH). Given an arbitrary view $\mathcal{I}$ defined by intrinsic and extrinsic parameters, we render the pixel at position $(u,v)$ with timestamp $t$ by blending visible Gaussians:
$\mathcal{I}(u,v,t)=$ $\sum_{i}f_{i}(u,v,t)\alpha_{i}\mathbf{c}_{i}(d)T_{i}$ (4)
with $\text{with }T_{i}=$ $\prod^{i-1}_{j=1}(1-f_{j}(u,v,t)\alpha_{j}).$ (5)
Here, $i$ indexes the visible Gaussians sorted by depth, $d$ refers to the direction of the pixel under the view $\mathcal{I}$, and $f_{i}(u,v,t)$ denotes the influence of a Gaussian on this position. To obtain the influence, unlike 3D Gaussian Splatting [12] which directly projects 3D Gaussians to image space, we need condition 4D Gaussians on time and then project them. More specifically, $f(u,v,t)$ is expressed by:
$f(u,v,t)=G_{t}(t|\mathbf{\mu}_{t},\Sigma_{t})G_{{P}(\mathbf{x})|t}(u,v|{P}(\mathbf{\mu}_{\mathbf{x}|t}),{P}(\Sigma_{\mathbf{x}|t})),$ (6)
where $G_{t}$ is the marginal distribution of the 4D Gaussian in time, and $G_{\text{P}(\mathbf{x})|t}$ is the projected version of the conditional 3D Gaussian with
$\mathbf{\mu}_{\mathbf{x}|t}$ $=\mathbf{\mu}_{\mathbf{x}}+\Sigma_{\mathbf{x},t}\Sigma_{t}^{-1}(t-\mathbf{\mu}_{t})$ (7)
$\Sigma_{\mathbf{x}|t}$ $=\Sigma_{\mathbf{x}}-\Sigma_{\mathbf{x},t}\Sigma_{t}^{-1}\Sigma_{t,\mathbf{x}}.$ (8)
The projection operation $P$ [12, 47] projects the world point $\mathbf{\mu}_{\mathbf{x}|t}$ to image space and transforms the covariance by ${P}(\Sigma_{\mathbf{x}|t})=JW\Sigma_{\mathbf{x}|t}W^{\top}J^{\top}$ where $W$ is the extrinsic matrix of $\mathcal{I}$ and $J$ is the Jacobian of the affine approximation of the projective transformation.
Optimization
During model training, optimization is performed on the mean ($\mathbf{\mu}$), covariance ($\Sigma$), opacity ($\alpha$), and spherical harmonic (SH) coefficients, as well as density control including densification and pruning for each Gaussian. The original objective function, as presented in [12], assumes clean training data, which may not be applicable to synthetic data with inherent imperfections. To address this, we introduce a confidence-aware objective loss formulation defined as:
$\mathcal{L}=\lambda_{\text{rgb}}\mathcal{C}_{\text{rgb}}\mathcal{L}_{\text{rgb}}+\lambda_{\text{ssim}}\mathcal{C}_{\text{ssim}}\mathcal{L}_{\text{ssim}},$ (9)
where $\mathcal{L}_{\text{rgb}}$ is the $L_{1}$ loss in RGB space, $\mathcal{L}_{\text{ssim}}$ is the SSIM loss [40], $\lambda_{\text{rgb/ssim}}$ is the respective weight hyper-parameter, and $\mathcal{C}_{\text{rgb/ssim}}$ is the confidence score of a generated image $\mathbf{I}$ calculated as
$\mathcal{C}_{\text{rgb}}=1-|\mathbf{I}-\hat{\mathbf{I}}|,\quad\mathcal{C}_{\text{ssim}}=\text{SSIM}(\mathbf{I},\hat{\mathbf{I}})$ (10)
where $\hat{\mathbf{I}}$ is the unwarped image from adjacent frames estimated by the optical flow method. In such way, the confidence $\mathcal{C}_{\text{rgb/ssim}}$ function as an adaptive role in controlling the loss and gradient by assigning lower weights to inconsistent regions.
4 Experiments
Input samples
Refer to caption Refer to caption Refer to caption Refer to caption
Consistent4D [10] ~165 min
Refer to caption Refer to caption Refer to caption Refer to caption
Ours images ~2 min
Refer to caption Refer to caption Refer to caption Refer to caption
Ours geometry ~12 min
Refer to caption Refer to caption Refer to caption Refer to caption
Figure 4: Qualitative comparisons. We show four images for each method, captured from two viewpoints at two timestamps.
4.1 Implementation details
Competitor
Video-to-4D is a relatively unexplored area, with only one existing method, Consistent4D [10]. In comparison, we obtained the results of Consistent4D by referencing its original paper and project page since the source code has not been publicly released.
Evaluation videos
To showcase the versatility of our proposed method, we conducted extensive experiments using a diverse set of data sources. Specifically, we focused on 20 sequences: alpaca, astronaut, bird, butterfly, cat, corgi, dinosaur, dog, dragon, duck, flower, guard, panda, penguin, rabbit, robot, snail, Spiderman, yellow face, yoxi.
For dragon, guard, and yoxi, 3D animated models were obtained from Sketchfab [2]. We rendered front-view videos and corresponding masks at 30 fps using Blender. The yellow face sequence was collected from the internet. To facilitate comparison with Consistent4D [10], we also conducted experiments on the videos used in their work [10].
Each sequence consists of monocular videos captured by a stationary camera, with durations ranging from approximately 1s to 6s. We uniformly selected 2-32 frames from each video as keyframes and segmented their foreground parts using SAM [13].
Table 1: Quantitative comparisons. The frame per second (FPS) is evaluated using a $256\times 256$ resolution rendering. ${}^{\dagger}$: Reported in [10] with one V100 GPU. ${}^{\ddagger}$: Measured on K-Planes [7], which partially serves as the fundamental 4D representation of Consistent4D [10].
Method CLIP-similarity $\uparrow$ Generation time $\downarrow$ FPS $\uparrow$ Explicit geometry
Consistent4D [10] 0.8438 165 ${}^{\dagger}$${}^{\ddagger}$
Efficient4D 0.8601 14 min ~1000
Implementation
For time-synchronous spatial volume, we typically set the weight as $\mathbf{w}=(0.1,0.1,0.6,0.1,0.1)$. In the reconstruction stage, we implement a modified CUDA rasterizer tailored to 4D Gaussian. Gaussians are initialized randomly within the space $[-0.6,0.6]^{3}$ with identity rotations and a initial number of 30,000-50,000. Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 9,000 iterations with a batch size of 4. Gaussians are densified every 50 iterations, with densification halted at the 7,000th iteration. All other hyperparameters remain consistent with those in [12]. For speed efficiency, our proposed method only costs about 2 minutes for image generation and 12 minutes for reconstruction on one A6000 GPU.
4.2 Qualitative evaluation
We present qualitative comparisons in Fig. 4. In geometry rendering, we extract points at the mean of Gaussians. We also eliminate white background points and transparent points with low opacity. When assessing texture quality, it is important to note that both methods fall under the category of lifting 2D to 4D. However, Consistent4D produces watercolor-like images with low fidelity, which can be attributed to conflicts from multiple supervisory signals during prolonged optimization. In contrast, our method excels in directly generating high-quality 2D images. Subsequently, our method’s reconstruction stage yields explicit point clouds and exhibits less degradation in rendering quality compared to the SDS-based Consistent4D. Overall, our method consistently outperforms Consistent4D in most cases.
4.3 Quantitative evaluation
In Tab. 1, we compare Consistent4D and our method based on average generation time, rendering time, and CLIP-similarity scores [31] for input image sequences and rendered images. Measurements used the image sequences described in Sec. 4.1. Consistent4D’s CLIP-similarity evaluation is limited to RGB images from 2-3 selected views in their paper and project page, while our evaluation includes 320 images uniformly distributed in space and time, covering 16 viewpoints and 20 timestamps, making our CLIP-similarity score more robust.
We draw several observation: (1) Our method achieves superior image quality. (2) Our method significantly accelerates the generation process, achieving over 10$\times$ speed improvement. (3) Our efficient 4D Gaussian representation and splatting based rendering enable real-time rendering and provide explicit point cloud geometry.
More specifically, the speed improvement is attributed to (1) our SDS-free design, converging optimization in fewer iterations, and (2) each iteration in our method requiring less time compared to Consistent4D. This time efficiency is further enhanced by avoiding recurrent evaluation in the diffusion model and prolonged forward rendering of the inefficient 4D representation as needed in the competitor at every iteration.
Refer to caption Refer to caption Refer to caption
Refer to caption Refer to caption Refer to caption
Input Generated images Geometry
Figure 5: Given only two input frames, our method is able to generate point clouds and temporal continuous images from novel views.
4.4 Sparse input evaluation
We assessed our method’s performance in an extremely sparse input scenario, comprising only two discrete frames. In contrast, Consistent4D fails to operate effectively under such conditions. In such cases, we set $w=0.25$ in Eq. 2. As illustrated in Fig. 5, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency.
Consider a scenario where we seek 4D modeling for a static toy. While the toy can take different poses, it lacks autonomous movement, posing a challenge for continuous video capturing. In these instances, our method demonstrates effectiveness by requiring only a few key frames to produce dynamic content, thereby expanding the potential applications of the video-to-4D task.
4.5 Ablation studies
Single-view
Refer to caption Refer to caption Refer to caption Refer to caption
No time-sync
Refer to caption Refer to caption Refer to caption Refer to caption
No interp
Refer to caption Refer to caption Refer to caption Refer to caption
Full setting
Refer to caption Refer to caption Refer to caption Refer to caption
Generated images Reconstructed Generated images Reconstructed
Figure 6: Ablation study on multi-view images generation, time-synchronous spatial volumes and frame interpolation. The images follow a chronological order from left to right.
Generated images
K-Plane [7] ~2 hrs
No conf ~12 min
Full setting ~12 min
Reconstructed novel views
Figure 7: Ablation study on 4D scene representation and confidence maps. The images follow a chronological order from left to right.
We performed ablation studies to assess the influence of different components in our approach. We compared our full method against five baseline settings: (1) Only input video are used for reconstruction; (2) Time-synchronous spatial volume is excluded; (3) Frame interpolation is excluded; (4) Confidence maps are excluded; (5) 4D Gaussian splatting is replaced with K-Plane [7]. The results corresponding to these settings can be found in Fig. 6 and Fig. 7.
Importance of synthetic data.
We compare our generated image matrix with utilizing only input video using the same proposed 4D Gaussian representation model. As shown in the first row of Fig. 6, when only relying on a single-view video, the model cannot produce any meaningful results for novel views. This indicates the importance of constructing proper training data.
Effect of time-synchronous spatial volume.
We assess the impact of the time-synchronous spatial volume concept introduced in SyncDreamer, as depicted in the contrast of the second and fourth rows in Fig. 6. Without time-synchronous spatial volume, the back aspects of the toy Spiderman exhibit inconsistencies, leading to distorted geometry. In contrast, the proposed time-synchronous spatial volume enhances both spatial and temporal consistency while preventing geometry collapse, resulting in more visually appealing image generation.
Effect of frame interpolation.
As illustrated in the contrast between the third and fourth rows in Fig. 6, frame interpolation is effective in mitigating the blurring observed in the rendered image from novel views. This is attributed to the low frame rates of the image matrix, which results in noticeable discontinuities.
Effect of confidence-aware in objective.
The integration of a confidence-aware loss in our design serves as a strategy to mitigate training data noise. This is evident in the contrast between the third and fourth rows in Fig. 7. The inclusion of confidence maps effectively reduces blurry rendering caused by inter-frame inconsistency, resulting in a significant overall improvement in quality.
Integrating existing 4D representation model.
Our synthetic training data is versatile and supports the optimization of 4D representation models, such as K-Plane [7]. In the comparison shown in Fig. 7, our 4D Gaussian splatting and K-Plane exhibit comparable proficiency in reconstructing dynamic scenes. Notably, our method surpasses K-Plane in speed, achieving a 10$\times$ improvement. Additionally, as emphasized in Sec. 4.3, our model enables real-time rendering with explicit point cloud geometry.
5 Conclusion
This study introduces a new framework, Efficient4D, designed for generating efficiently dynamic 4D objects seamlessly from monocular videos captured by a stationary camera. The Efficient4D consists of two main stages: first, generating consistent multi-view videos with spatial and temporal coherence, and second, rapidly producing 4D object reconstructions using an innovative 4D Gaussian splatting model. Our approach, utilizing SDS-Free design and efficient 4D representation, significantly accelerates the generation process, achieving over 10 times faster speeds compared to previous work, while still delivering superior reconstruction and novel view synthesis results. Moreover, our model is effective in extremely sparse input scenarios, requiring only two available images, thereby expanding its application scope.
Limitations.
Our local smoothing approach in sliding window style with our image sequence generator poses challenges in handling long-duration videos. To address this issue, replacing the smoothing filtering layer with a learnable attention layer featuring global receptive fields is a viable solution, particularly when dealing with extensive 4D datasets. Additionally, handling long videos may necessitate significant GPU memory, which can be mitigated by employing multi-GPUs or CPUs, albeit with an increase in processing time.

### v2

@@
-Fast Dynamic 3D Object Generation from a Single-view Video
+https://fudan-zvg.github.io/Efficient4D
 Abstract
-Generating dynamic three-dimensional (3D) object from a single-view video is challenging due to the lack of 4D labeled data.
-Existing methods extend text-to-3D pipelines by transferring off-the-shelf image generation models such as score distillation sampling, but they are slow and expensive to scale (e.g., 150 minutes per object) due to the need for back-propagating the information-limited supervision signals through a large pretrained model.
-To address this limitation, we propose an efficient video-to-4D object generation framework called Efficient4D.
+Generating dynamic 3D object from a single-view video is challenging due to the lack of 4D labeled data.
+Extending image-to-3D pipelines by transferring off-the-shelf image generation models such as score distillation sampling, existing methods tend to be slow and expensive to scale due to the need for back-propagating the information-limited supervision signals through a large pretrained model.
+To address this, we propose an efficient video-to-4D object generation framework called Efficient4D.
 It generates high-quality spacetime-consistent images under different camera views, and then uses them as labeled data to directly train a novel 4D Gaussian splatting model with explicit point cloud geometry, enabling real-time rendering under continuous camera trajectories.
-Extensive experiments on synthetic and real videos show that Efficient4D offers a remarkable 10-fold increase in speed when compared to prior art alternatives while preserving the same level of innovative view synthesis quality.
-For example, Efficient4D takes only 14 minutes to model a dynamic object.
-11footnotetext:  Corresponding to lizhangfd@fudan.edu.cn
+Extensive experiments on synthetic and real videos show that Efficient4D offers a remarkable 20-fold increase in speed when compared to prior art alternatives while preserving the quality of novel view synthesis.
+For example, Efficient4D takes only 6 mins to model a dynamic object, vs 120 mins by Consistent4D.
+Keywords:
+4D Generation Gaussian Splatting Video Efficiency
 1 Introduction
 Humans possess a remarkable capacity to comprehensively comprehend the spatial and temporal characteristics of a dynamic object in a brief video, even with a limited perspective, enabling them to predict its appearance in unseen viewpoints over time.
-Despite the significant advancement of three-dimensional (3D) object generation, existing works [28, 17, 4, 41] mostly consider static scenes or objects.
+Despite the significant advancement of 3D object generation, existing works [29, 18, 4, 43] mostly consider static scenes or objects.
 With the availability of large-scale 3D datasets [6, 5], training generalizable models capable of directly generating multi-view images becomes possible [21, 22, 23].
@@
 By further augmenting these generated static objects with predefined animations [1], dynamic 3D content can be generated.
-However, this is highly limited due to the need for fine-grained meshes as well as rigid restrictions.
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Input Generated dynamic objects
+However, this approach is still limited due to the need for fine-grained meshes as well as rigid restrictions.
+Refer to caption  Refer to caption Refer to caption
+ Refer to caption  Refer to caption Refer to caption
+Input Consistent4D [13] Efficient4D(Ours)
 Figure 1: Examples of video-to-4D generation. Input: A brief video of a dynamic object, as represented by 3 frames per case; Output: Generated novel views at different timestamps.
@@
 In quest of the aforementioned human’s capability, a recent work [13] proposed to generate dynamic 3D object images from a single-view video (statically captured monocular video from a fixed view), namely as video-to-4D object generation.
-However, similar as [36] this method is also slow to train (e.g., 150 minutes to model a single dynamic object) in addition to complex design, hence unscalable in practice.
-To address identified limitations, we formulate an efficient video-to-4D object generation method in a two-staged pipeline called Efficient4D.
+However, similar as [38] this method is also slow to train (e.g., 150 minutes to model a single dynamic object) in addition to complex design, hence unscalable and expensive in practice.
+To address identified limitations, we formulate an efficient video-to-4D two staged object generation method called Efficient4D.
 In the first stage, we generate spacetime-consistent images across different camera views as synthetic training data.
-This is realized by imposing temporal smoothing filter into a previous multi-view image generator (e.g., SyncDreamer [21]) in tandem with frame interpolation.
+This is realized by imposing temporal smoothing into a multi-view image generator (e.g., SyncDreamer [22]) in tandem with frame interpolation.
 In the second stage, we use these training data to optimize a novel 4D Gaussian splatting model.
-This is an extension of the recent 3D Gaussian splatting [12], originally designed for static 3D scene representation, with the temporal dimension introduced additionally on top.
-Using Gaussian representation brings about further computational efficiency gain along with the explicit geometry property, when compared with NeRF based designs.
-To tackle the discontinuity between generated frames, we design a inconsistency-aware loss function.
-A couple of video-to-4D results are presented in Fig.
-1.
+This is an extension of the 3D Gaussian splatting [15], originally designed for static 3D scene representation, with the temporal dimension introduced additionally.
+Using Gaussian representation brings about further computational efficiency gain along with the explicit geometry property, when compared with NeRF based designs (Fig.
+1).
+To tackle the challenging discontinuity between generated frames, we design an inconsistency-aware loss function.
 Our contributions are summarized as follows: (i) We consider for the first time the efficiency challenge with the under-studied video-to-4D object generation problem.
 (ii) We propose an efficient video-to-4D object generation pipeline, Efficient4D, characterized by directly generating high-quality training data without the need for heavy supervision back-propagation through a large pretrained model as suffered by most 3D/4D object generation approaches.
-(iii) Instead of using NeRF as the 3D representation model, we introduce a novel 4D Guardian splatting model with the real-time rendering advantage.
-(iv) Extensive experiments on both synthetic and real videos validate the significant efficiency advantage (e.g., 10$\times$ speedup) of our Efficient4D over the prior art, whilst maintaining the same level of new view synthesis.
-Additionally, our method can also work well under the more challenging few-shot setting where only a handful of key frames are available for training, further extending the application scope.
+(iii) Instead of using NeRF as the 3D representation model, we introduce a novel 4D Gaussian splatting model with the real-time rendering advantage.
+(iv) Extensive experiments on both synthetic and real videos validate the significant efficiency advantage (e.g., 20$\times$ speedup) of our Efficient4D over the prior art, whilst maintaining the quality of novel view synthesis.
+Also, our method can work well under the more challenging few-shot setting where only a handful of key frames are available for training, further extending the application scope.
 2 Related work
-4D representation
-Efforts to synthesize videos with free viewpoint control in dynamic scenes have a well-documented history [46].
-For example, pre-NeRF [25] approaches faced challenges in reconstructing intricate scene details.
-Recent advancements in 4D representations, particularly those based on neural rendering, include D-NeRF [29], which decouples geometry and motion, utilizing a canonical space and a learned deformation field.
-HyperNeRF [27] extends the deformation field to a higher dimension, employing the level-set method for smoother optimization.
-DeVRF [18] reconstructs the canonical space from multi-view dense observations in the initial frame and learns the deformation field with sparse views in subsequent timesteps, utilizing a voxel grid for representation.
-Similarly, deformable 3D Gaussians [44] integrates the deformation field with 3D Gaussian Splatting for the scene’s geometric structure.
-DyNeRF [14] extends NeRF with an additional latent code to capture detailed temporal variations, while DynIBaR [16] extends the image-based rendering paradigm for representing long videos with complex camera and object motions.
-Another group of method [7, 3, 33] adopts tensor decomposition of 4D volumes to represent dynamic 3D scenes.
-However, these representations often involve complex training schedules or suffer from slow convergence.
-In this work, drawing inspiration from 3D Gaussian Splatting [12], we propose an efficient 4D scene primitive for the representation of dynamic scenes.
 3D generation
@@
 Importantly, SyncDreamer [22] produces multi-view consistent images, offering inspiration for reconstructing 3D objects.
+4D representation
+Efforts to synthesize videos with free-viewpoint control in dynamic scenes have a well-documented history [53].
+For example, pre-NeRF [26] approaches challenges in reconstructing intricate scene details.
+Recent advancements in 4D representations, particularly those based on neural rendering, include D-NeRF [30], DeVRF [19] and HyperNeRF [28], which decouple geometry and motion, utilizing a canonical space and a learned deformation field.
+DynIBaR [17] deploys an image-based rendering paradigm for representing long videos with complex camera and object motions.
+Another group of methods [8, 3, 35] adopt tensor decomposition of 4D volumes to represent dynamic 3D scenes.
+Recently, Gaussian Splatting [15] has received widespread attention for its real-time high-fidelity rendering, especially its explicit point-based representation which holds great potential in modeling dynamic scenes.
+Consequently, a significant amount of work has been proposed to explore its extension to dynamic scene modeling.
+Among them, Deformable 3D Gaussians [49] and 4DGaussian [46] integrated the deformation field with 3D Gaussian Splatting for the joint learning of the scene’s geometry and dynamics.
+SC-GS [10] represents the motion with a set of sparse control points to achieve reconstruction and motion editing.
+Unlike the previous radiance field-based representations often involving complex training schedules or suffering from slow convergence, these methods can be optimized efficiently, achieving real-time rendering while surpassing the past methods in terms of quality.
+Our work is perpendicular to all the above works, where any of them can be deployed in our reconstruction phase.
+But considering both optimization efficiency and expressive capability, we propose to represent dynamic 3D assets by a set of native 4D scene primitives.
 4D generation
@@
 Instead of text input, Consistent4D [13] conditions the generation of 4D object over time on a monocular video with richer and more specific information.
-However, this work is computationally inefficient due to inheriting the previous SDS loss, along with complex pipeline design.
-To overcome this limitation, we present a novel two-staged pipeline in a generation-and-reconstruction strategy, drastically boosting the model training speed by 10$\times$ whilst maintaining the quality of novel view synthesis.
+However, it is computationally inefficient due to inheriting the previous SDS loss, along with complex pipeline design.
+To overcome this limitation, we present a novel two-staged pipeline in a generation-and-reconstruction strategy, drastically boosting the training speed by 20$\times$ whilst maintaining the quality of novel view synthesis.
+3 Method
 Figure 2: Overview of our Efficient4D approach.
-Given as the input a brief video depicting a dynamic object in limited perspectives, our model aims to generate this object with geometrical and temporal consistency under any specific view and time.
+Given as the input a brief video depicting a dynamic object from a single perspectives, our model aims to generate this object with geometrical and temporal consistency under any specific view and time.
 Efficient4D comprises two components: (A) Image sequence synthesis across views and timestamps, resulting in (c) an image matrix where each row consists of multi-view geometrically consistent images and each column consists of view-specific temporally consistent images.
@@
 It can be trained efficiently and robustly under the confidence-aware (e) supervision on the generated image matrix.
-3 Method
-Our Efficient4D addresses the challenge of efficiently generating dynamic objects under novel views from a single-view video . The model, illustrated in Fig. 2, comprises two key components:
+Our Efficient4D addresses the challenge of efficiently generating dynamic objects under novel views from a single-view video. As illustrated in Fig. 2, it comprises two key components:
   • An image synthesis pipeline (Sec. 3.1) generates images across views and timestamps, ensuring sufficient geometry and temporal consistency.
@@
 This synchronization ensures consistent features across $k$ past and $k$ future frames during the denoising process, enhancing temporal consistency.
+With this time-synchronized spatial volumes, the proposed SyncDreamer-T is entirely training-free established on the pretrained SyncDreamer.
 To further improve temporal resolution, video frame interpolation (e.g., RIFE [12]) is applied after generating the image matrix $\mathcal{D}$ as described above.
@@
 This results in a total of $M=4K-3$ images in a column of $\mathcal{D}$.
+Refer to caption Refer to caption
+Input Feature fusion
+Figure 3: Example analysis on temporal synchronization.
+In this illustration, we manipulate the fusion ratio ($w$) across a range from 0 to 0.5 for the spatial feature volumes of two input frames.
+The results indicate that a moderate ratio can achieve superior outcomes by balancing both temporal consistency and motion independence.
 Analysis on temporal synchronization
@@
 $\tilde{\mathcal{V}}_{i}=(1-w)\mathcal{V}_{i}+w\mathcal{V}_{3-i},$ (2)
-where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion” as illustrated in Fig.
+where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion" as illustrated in Fig.
 3.
@@
 In practical terms, it is recommended to set the ratio between 0.25 and 0.4, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry.
-Refer to caption Refer to caption
-Input Feature fusion
-Figure 3: Example analysis on temporal synchronization.
-In this illustration, we manipulate the fusion ratio ($w$) across a range from 0 to 0.5 for the spatial feature volumes of two input frames.
-The results indicate that a moderate ratio can achieve superior outcomes by balancing both temporal consistency and motion independence.
+Note that the results in Fig.
+3 are not sensitive to the choice in this range.
 3.2 4D generation through reconstruction
-Recall that we desire a 3D dynamic object, we should not be content with the discrete images.
+Aiming for 3D dynamic object modeling, discrete images not suffice.
 Our next goal is to model a truly 4D content from the image matrix $\mathcal{D}$.
@@
 where $G_{t}$ is the marginal distribution of the 4D Gaussian in time, and $G_{\text{P}(\mathbf{x})|t}$ is the projected version of the conditional 3D Gaussian with
-$\mathbf{\mu}_{\mathbf{x}|t}$ $=\mathbf{\mu}_{\mathbf{x}}+\Sigma_{\mathbf{x},t}\Sigma_{t}^{-1}(t-\mathbf{\mu}_{t})$ (7)
+$\mathbf{\mu}_{\mathbf{x}|t}$ $=\mathbf{\mu}_{\mathbf{x}}+\Sigma_{\mathbf{x},t}\Sigma_{t}^{-1}(t-\mathbf{\mu}_{t}),$ (7)
 $\Sigma_{\mathbf{x}|t}$ $=\Sigma_{\mathbf{x}}-\Sigma_{\mathbf{x},t}\Sigma_{t}^{-1}\Sigma_{t,\mathbf{x}}.$ (8)
@@
 Optimization
-During model training, optimization is performed on the mean ($\mathbf{\mu}$), covariance ($\Sigma$), opacity ($\alpha$), and spherical harmonic (SH) coefficients, as well as density control including densification and pruning for each Gaussian.
-The original objective function, as presented in [12], assumes clean training data, which may not be applicable to synthetic data with inherent imperfections.
+In training, optimization is performed on the mean ($\mathbf{\mu}$), covariance ($\Sigma$), opacity ($\alpha$), and spherical harmonic (SH) coefficients, as well as density control including densification and pruning for each Gaussian.
+The original objective function, as presented in [15], assumes clean training data, which may not less valid for synthetic data with inherent imperfections.
 To address this, we introduce a confidence-aware objective loss formulation defined as:
@@
 $\mathcal{C}_{\text{rgb}}=1-|\mathbf{I}-\hat{\mathbf{I}}|,\quad\mathcal{C}_{\text{ssim}}=\text{SSIM}(\mathbf{I},\hat{\mathbf{I}})$ (10)
-where $\hat{\mathbf{I}}$ is the unwarped image from adjacent frames estimated by the optical flow method.
+where $\hat{\mathbf{I}}$ is the unwarped image from adjacent frames estimated by optical flow.
 In such way, the confidence $\mathcal{C}_{\text{rgb/ssim}}$ function as an adaptive role in controlling the loss and gradient by assigning lower weights to inconsistent regions.
 4 Experiments
+4.1 Implementation details
+C4D [13]
+Ours
+GT
+Figure 4: Qualitative evaluation of Consistent4D (C4D) [13] and our method against ground truth (GT) on synthetic data.
+Competitors
+Video-to-4D is a relatively unexplored area, with only one existing method, Consistent4D [13].
+In comparison, we obtained the results of Consistent4D by running its released code.
+We also compare SyncDreamer [22] partially by replacing SyncDreamer-T in Sec.
+4.5.
+Evaluation videos
+To showcase the versatility of our proposed method, we conducted extensive experiments using a diverse set of data sources.
+Specifically, we focused on 36 sequences: 32 sequences released by [13] and 4 sequences processed by ourselves.
+Among the released data, seven sequences are synthetic data where ground truth are provided.
+Our four sequences are used for sparse input evaluation in Sec.
+4.4, which only contain two frames for each sequence.
+Three of them, named dragon, guard, and yoxi, are rendered from 3D animated models obtained from Sketchfab [2].
+The other one named yellow face was collected from the internet.
+Each sequence consists of monocular videos captured by a stationary camera, with duration ranging from approximately 1s to 3s.
+Evaluation metrics
+As 4D generation research is still at early stage, there is no well established metric yet.
+However, we adopt multiple metrics by referring related works for comprehensive evaluation.
+For evaluating on synthetic data, we use LPIPS score [51] and CLIP similarity [32] between rendered images and ground truth following [13].
+For the cases without ground truth, we also use CLIP-similarity between generated images and input frames as [39, 33] to measure image quality.
+For temporal smoothness, following [50, 7, 9] we use CLIP-T to measure the similarity between adjacent frames of a generated video from different views, including front (CLIP-T-f), side (CLIP-T-s) and back (CLIP-T-b) views.
+To evaluate a 4D object completely for different methods, we obtain the generated images by rendering 320 images uniformly distributed in space and time, covering 16 viewpoints and 20 timestamps.
+Implementation
+For SyncDreamer-T, we typically set the weight as $\mathbf{w}=(0.1,0.1,0.6,0.1,0.1)$ in time-synchronous spatial volume.
+The improved sampling strategy [45] can be seamlessly integrated into SyncDreamer-T.
+In the reconstruction stage, we implement a modified CUDA rasterizer tailored to 4D Gaussian.
+Gaussians are initialized randomly within the space $[-0.6,0.6]^{3}$ with identity rotations and a initial number of 20,000.
+Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 20,000 iterations with a batch size of 2.
+Gaussians are densified every 200 iterations, with densification halted at the 10,000th iteration.
+All other hyperparameters remain consistent with those in [15].
+For speed efficiency, our proposed method only costs about 2 minutes for image generation and 4 minutes for reconstruction on one A6000 GPU.
+4.2 Qualitative evaluation
 Input samples
@@
 Refer to caption Refer to caption Refer to caption Refer to caption
-Ours images ~2 min
+Our images ~2 min
 Refer to caption Refer to caption Refer to caption Refer to caption
-Ours geometry ~12 min
+Our geometry ~4 min
 Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 4: Qualitative comparisons. We show four images for each method, captured from two viewpoints at two timestamps.
-4.1 Implementation details
-Competitor
-Video-to-4D is a relatively unexplored area, with only one existing method, Consistent4D [10].
-In comparison, we obtained the results of Consistent4D by referencing its original paper and project page since the source code has not been publicly released.
-Evaluation videos
-To showcase the versatility of our proposed method, we conducted extensive experiments using a diverse set of data sources.
-Specifically, we focused on 20 sequences: alpaca, astronaut, bird, butterfly, cat, corgi, dinosaur, dog, dragon, duck, flower, guard, panda, penguin, rabbit, robot, snail, Spiderman, yellow face, yoxi.
-For dragon, guard, and yoxi, 3D animated models were obtained from Sketchfab [2].
-We rendered front-view videos and corresponding masks at 30 fps using Blender.
-The yellow face sequence was collected from the internet.
-To facilitate comparison with Consistent4D [10], we also conducted experiments on the videos used in their work [10].
-Each sequence consists of monocular videos captured by a stationary camera, with durations ranging from approximately 1s to 6s.
-We uniformly selected 2-32 frames from each video as keyframes and segmented their foreground parts using SAM [13].
-Table 1: Quantitative comparisons.
-The frame per second (FPS) is evaluated using a $256\times 256$ resolution rendering.
-${}^{\dagger}$: Reported in [10] with one V100 GPU.
-${}^{\ddagger}$: Measured on K-Planes [7], which partially serves as the fundamental 4D representation of Consistent4D [10].
-Method CLIP-similarity $\uparrow$ Generation time $\downarrow$ FPS $\uparrow$ Explicit geometry
-Consistent4D [10] 0.8438 165 ${}^{\dagger}$${}^{\ddagger}$
-Efficient4D 0.8601 14 min ~1000
-Implementation
-For time-synchronous spatial volume, we typically set the weight as $\mathbf{w}=(0.1,0.1,0.6,0.1,0.1)$.
-In the reconstruction stage, we implement a modified CUDA rasterizer tailored to 4D Gaussian.
-Gaussians are initialized randomly within the space $[-0.6,0.6]^{3}$ with identity rotations and a initial number of 30,000-50,000.
-Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 9,000 iterations with a batch size of 4.
-Gaussians are densified every 50 iterations, with densification halted at the 7,000th iteration.
-All other hyperparameters remain consistent with those in [12].
-For speed efficiency, our proposed method only costs about 2 minutes for image generation and 12 minutes for reconstruction on one A6000 GPU.
-4.2 Qualitative evaluation
-We present qualitative comparisons in Fig.
-4.
+Figure 5: Qualitative comparisons. We show four images per method, captured from two viewpoints at two timestamps. Our geometry is the points at 3D mean of Gaussians.
+We first present the results on synthetic data in Fig.
+4 with ground truth shown.
+Our visual results exhibit superior accuracy in both texture and geometry when compared to the ground truth.
+For other data without ground truth, we present qualitative comparisons in Fig.
+5.
 In geometry rendering, we extract points at the mean of Gaussians.
@@
 4.3 Quantitative evaluation
-In Tab.
-1, we compare Consistent4D and our method based on average generation time, rendering time, and CLIP-similarity scores [31] for input image sequences and rendered images.
-Measurements used the image sequences described in Sec.
-4.1.
-Consistent4D’s CLIP-similarity evaluation is limited to RGB images from 2-3 selected views in their paper and project page, while our evaluation includes 320 images uniformly distributed in space and time, covering 16 viewpoints and 20 timestamps, making our CLIP-similarity score more robust.
-We draw several observation: (1) Our method achieves superior image quality.
-(2) Our method significantly accelerates the generation process, achieving over 10$\times$ speed improvement.
+Table 1: Quantitative comparisons. The frame per second (FPS) is evaluated using a $256\times 256$ resolution rendering. : Measured using code released by Consistent4D [13] with one A6000 GPU.
+Consistent4D [13] Efficient4D(Ours)
+- Comparison with ground truth on synthetic data
+CLIP $\uparrow$ 0.87 0.92
+LPIPS $\downarrow$ 0.16 0.14
+- Comparison on all cases
+CLIP $\uparrow$ 0.8471 0.8721
+CLIP-T-f $\uparrow$ 0.9692 0.9702
+CLIP-T-s $\uparrow$ 0.9658 0.9726
+CLIP-T-b $\uparrow$ 0.9697 0.9707
+Generation time $\downarrow$ 120 mins 6 mins
+FPS $\uparrow$ ~20 ~1000
+Explicit geometry
+In Tab. 1, we compare Consistent4D and our method quantitatively. The metrics and data used are described in Sec. 4.1.
+We draw several observations: (1) For synthetic data, our generation results produce better CLIP and LPIPS scores.
+(2) Our method achieves superior image quality and temporal smoothness on all cases.
+(3) Our method significantly accelerates the generation process, achieving over 20$\times$ speed improvement.
 (4) Our efficient 4D Gaussian representation and splatting based rendering enable real-time rendering and provide explicit point cloud geometry.
-More specifically, the speed improvement is attributed to (1) our SDS-free design, converging optimization in fewer iterations, and (2) each iteration in our method requiring less time compared to Consistent4D.
+More specifically, the speed improvement is attributed to (1) our SDS-free design, converging optimization faster, and (2) each iteration in our method requiring much less time compared to Consistent4D.
 This time efficiency is further enhanced by avoiding recurrent evaluation in the diffusion model and prolonged forward rendering of the inefficient 4D representation as needed in the competitor at every iteration.
@@
 Refer to caption Refer to caption Refer to caption
-Input Generated images Geometry
-Figure 5: Given only two input frames, our method is able to generate point clouds and temporal continuous images from novel views.
+Input Ours Consistent4D [13]
+Figure 6: Given only two input frames, our method is able to generate smooth dynamics. For each case, we show three internal images from two novel views.
 4.4 Sparse input evaluation
 We assessed our method’s performance in an extremely sparse input scenario, comprising only two discrete frames.
-In contrast, Consistent4D fails to operate effectively under such conditions.
 In such cases, we set $w=0.25$ in Eq.
@@
 As illustrated in Fig.
-5, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency.
+6, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency.
+In contrast, Consistent4D fails to operate effectively under such conditions.
 Consider a scenario where we seek 4D modeling for a static toy.
@@
 4.5 Ablation studies
-Single-view
+Single view
 Refer to caption Refer to caption Refer to caption Refer to caption
@@
 Refer to caption Refer to caption Refer to caption Refer to caption
-Generated images Reconstructed Generated images Reconstructed
-Figure 6: Ablation study on multi-view images generation, time-synchronous spatial volumes and frame interpolation. The images follow a chronological order from left to right.
-Generated images
-K-Plane [7] ~2 hrs
-No conf ~12 min
-Full setting ~12 min
+Input/Generated images Rendered Input/Generated images Rendered
+Figure 7: Ablation study on image generation, time-synchronous spatial volumes and frame interpolation. The images follow a chronological order from left to right.
+K-Plane [8] PSNR: 33.50 Time: 2 hrs
+Refer to caption Refer to caption
+No conf PSNR: 32.13 Time: 6 mins
+Refer to caption Refer to caption
+Full setting PSNR: 32.64 Time: 6 mins
+Refer to caption Refer to caption
 Reconstructed novel views
 Figure 8: Ablation study on 4D scene representation and confidence maps. The images follow a chronological order from left to right.
+Table 2: Quantitative ablation study on all our modules mesured on 4 sequences: alpaca, astronaut, rabbit and spiderman.
+Method CLIP $\uparrow$ CLIP-T-f $\uparrow$ CLIP-T-s $\uparrow$ CLIP-T-b $\uparrow$
+Single-view 0.6684 - - -
+No time-sync 0.8270 0.9258 0.9197 0.8819
+No interp 0.8595 0.9409 0.9443 0.9336
+No conf 0.8611 0.9658 0.9592 0.9611
+Full setting 0.8702 0.9689 0.9673 0.9684
 We performed ablation studies to assess the influence of different components in our approach.
 We compared our full method against five baseline settings: (1) Only input video are used for reconstruction; (2) Time-synchronous spatial volume is excluded; (3) Frame interpolation is excluded; (4) Confidence maps are excluded; (5) 4D Gaussian splatting is replaced with K-Plane [8].
+Note setting (2) is equivalent to the case where SyncDreamer-T is replaced with original SyncDreamer.
 The results corresponding to these settings can be found in Fig.
 7 and Fig.
-7.
+8.
+Table 2 also gives quantitative evaluations.
 Importance of synthetic data.
@@
 As shown in the first row of Fig.
-6, when only relying on a single-view video, the model cannot produce any meaningful results for novel views.
+7, when only relying on a single-view video, the model cannot produce any meaningful results for novel views.
 This indicates the importance of constructing proper training data.
 Effect of time-synchronous spatial volume.
-We assess the impact of the time-synchronous spatial volume concept introduced in SyncDreamer, as depicted in the contrast of the second and fourth rows in Fig.
-6.
+We assess the impact of the time-synchronous spatial volume concept introduced in SyncDreamer, as depicted in the contrast of the second and last rows in Fig.
+7 and Tab.
+2.
 Without time-synchronous spatial volume, the back aspects of the toy Spiderman exhibit inconsistencies, leading to distorted geometry.
-In contrast, the proposed time-synchronous spatial volume enhances both spatial and temporal consistency while preventing geometry collapse, resulting in more visually appealing image generation.
+In contrast, the proposed time-synchronous spatial volume enhances both spatial and temporal consistency while preventing geometry collapse, resulting in more visually appealing image generation and higher CLIP-T scores.
 Effect of frame interpolation.
-As illustrated in the contrast between the third and fourth rows in Fig.
-6, frame interpolation is effective in mitigating the blurring observed in the rendered image from novel views.
+As illustrated in the contrast between the third and last rows in Fig.
+7 and Tab.
+2, frame interpolation is effective in mitigating the blurring observed in the rendered image from novel views, thus delivering higher CLIP-T scores.
 This is attributed to the low frame rates of the image matrix, which results in noticeable discontinuities.
@@
 The integration of a confidence-aware loss in our design serves as a strategy to mitigate training data noise.
-This is evident in the contrast between the third and fourth rows in Fig.
-7.
-The inclusion of confidence maps effectively reduces blurry rendering caused by inter-frame inconsistency, resulting in a significant overall improvement in quality.
+This is evident in the contrast between the “No conf” and “Full setting” results in Fig.
+8 and Tab.
+2.
+The inclusion of confidence maps effectively reduces blurry rendering caused by inter-frame inconsistency, resulting in a significant overall improvement in quality and temporal smoothness.
 Integrating existing 4D representation model.
@@
 In the comparison shown in Fig.
-7, our 4D Gaussian splatting and K-Plane exhibit comparable proficiency in reconstructing dynamic scenes.
-Notably, our method surpasses K-Plane in speed, achieving a 10$\times$ improvement.
+8, our 4D Gaussian splatting and K-Plane exhibit comparable proficiency in reconstructing dynamic scenes.
+Notably, our method surpasses K-Plane in speed, achieving a 20$\times$ improvement.
 Additionally, as emphasized in Sec.
@@
 The Efficient4D consists of two main stages: first, generating consistent multi-view videos with spatial and temporal coherence, and second, rapidly producing 4D object reconstructions using an innovative 4D Gaussian splatting model.
-Our approach, utilizing SDS-Free design and efficient 4D representation, significantly accelerates the generation process, achieving over 10 times faster speeds compared to previous work, while still delivering superior reconstruction and novel view synthesis results.
+Our approach, utilizing SDS-Free design and efficient 4D representation, significantly accelerates the generation process, achieving over 20 times faster speeds compared to previous work, while still delivering superior reconstruction and novel view synthesis results.
 Moreover, our model is effective in extremely sparse input scenarios, requiring only two available images, thereby expanding its application scope.

### v3

@@
-https://fudan-zvg.github.io/Efficient4D
+Efficient4D: Fast Dynamic 3D Object Generation from a Single-view Video
 Abstract
 Generating dynamic 3D object from a single-view video is challenging due to the lack of 4D labeled data.
-Extending image-to-3D pipelines by transferring off-the-shelf image generation models such as score distillation sampling, existing methods tend to be slow and expensive to scale due to the need for back-propagating the information-limited supervision signals through a large pretrained model.
+An intuitive approach is to extend previous image-to-3D pipelines by transferring off-the-shelf image generation models such as score distillation sampling.
+However, this approach would be slow and expensive to scale due to the need for back-propagating the information-limited supervision signals through a large pretrained model.
 To address this, we propose an efficient video-to-4D object generation framework called Efficient4D.
-It generates high-quality spacetime-consistent images under different camera views, and then uses them as labeled data to directly train a novel 4D Gaussian splatting model with explicit point cloud geometry, enabling real-time rendering under continuous camera trajectories.
-Extensive experiments on synthetic and real videos show that Efficient4D offers a remarkable 20-fold increase in speed when compared to prior art alternatives while preserving the quality of novel view synthesis.
-For example, Efficient4D takes only 6 mins to model a dynamic object, vs 120 mins by Consistent4D.
+It generates high-quality spacetime-consistent images under different camera views, and then uses them as labeled data to directly reconstruct the 4D content through a 4D Gaussian splatting model.
+Importantly, our method can achieve real-time rendering under continuous camera trajectories.
+To enable robust reconstruction under sparse views, we introduce inconsistency-aware confidence-weighted loss design, along with a lightly weighted score distillation loss.
+Extensive experiments on both synthetic and real videos show that Efficient4D offers a remarkable 10-fold increase in speed when compared to prior art alternatives while preserving the quality of novel view synthesis.
+For example, Efficient4D takes only 10 minutes to model a dynamic object, vs 120 minutes by the previous art model Consistent4D.
+Our code is publicly available at https://github.com/fudan-zvg/Efficient4D.
 keywords:
-4D Generation Gaussian Splatting Video Efficiency
+4D Generation, Efficiency, Diffusion Model, Gaussian Splatting, Video
 1 Introduction
@@
 However, this approach is still limited due to the need for fine-grained meshes as well as rigid restrictions.
-Refer to caption  Refer to caption Refer to caption
- Refer to caption  Refer to caption Refer to caption
-Input Consistent4D [13] Efficient4D(Ours)
-Figure 1: Examples of video-to-4D generation. Input: A brief video of a dynamic object, as represented by 3 frames per case; Output: Generated novel views at different timestamps.
 Directly generating 4D object/scene content from text description has been recently attempted [13].
@@
 In quest of the aforementioned human’s capability, a recent work [14] proposed to generate dynamic 3D object images from a single-view video (statically captured monocular video from a fixed view), namely as video-to-4D object generation.
-However, similar as [38] this method is also slow to train (e.g., 150 minutes to model a single dynamic object) in addition to complex design, hence unscalable and expensive in practice.
+However, similar as [13] this method is also slow to train (e.g., 120 minutes to model a single dynamic object) in addition to complex design, hence unscalable and expensive in practice.
 To address identified limitations, we formulate an efficient video-to-4D two staged object generation method called Efficient4D.
@@
 This is realized by imposing temporal smoothing into a multi-view image generator (e.g., SyncDreamer [8]) in tandem with frame interpolation.
-In the second stage, we use these training data to optimize a novel 4D Gaussian splatting model.
-This is an extension of the 3D Gaussian splatting [15], originally designed for static 3D scene representation, with the temporal dimension introduced additionally.
-Using Gaussian representation brings about further computational efficiency gain along with the explicit geometry property, when compared with NeRF based designs (Fig.
-1).
-To tackle the challenging discontinuity between generated frames, we design an inconsistency-aware loss function.
+In the second stage, we use these training data to optimize a 4D Gaussian splatting model [15].
+This is an extension of the 3D Gaussian splatting [16], originally designed for static 3D scene representation, with the temporal dimension introduced additionally, allowing for real-time rendering under continuous camera trajectories.
+Using Gaussian representation brings about further computational efficiency gain, when compared with NeRF based designs (Figure 1).
+To tackle the challenging discontinuity between the generated sparse frames, we design an inconsistency-aware loss function based on whether there is confidence in the consistency of a frame with its adjacent frames.
+Along with a lightly weighted score distillation sampling loss for smooth viewpoint transitions, this technique enables robust 4D reconstruction.
+Notably, although our method takes a single-view video as input, Efficient4Dcan be easily extended to image-to-4D task by leveraging a image-to-video diffusion model [17].
 Our contributions are summarized as follows: (i) We consider for the first time the efficiency challenge with the under-studied video-to-4D object generation problem.
 (ii) We propose an efficient video-to-4D object generation pipeline, Efficient4D, characterized by directly generating high-quality training data without the need for heavy supervision back-propagation through a large pretrained model as suffered by most 3D/4D object generation approaches.
-(iii) Instead of using NeRF as the 3D representation model, we introduce a novel 4D Gaussian splatting model with the real-time rendering advantage.
-(iv) Extensive experiments on both synthetic and real videos validate the significant efficiency advantage (e.g., 20$\times$ speedup) of our Efficient4D over the prior art, whilst maintaining the quality of novel view synthesis.
+We also extends Efficient4D to image-to-4D task.
+(iii) We introduce a inconsistency-aware confidence-weighted loss for reconstructing a 4D Gaussian splatting model using the generated training data efficiently and robustly.
+(iv) Extensive experiments on both synthetic and real data validate the significant efficiency advantage (e.g., 10$\times$ speedup) of our Efficient4D over the prior art, whilst maintaining the quality of novel view synthesis.
 Also, our method can work well under the more challenging few-shot setting where only a handful of key frames are available for training, further extending the application scope.
+Input Consistent4D [14]: 120 mins Efficient4D (Ours): 10 mins
+Input 4DGen [18]: 130 mins Efficient4D (Ours): 10 mins
+Input STAG4D [19]: 70 mins Efficient4D (Ours): 10 mins
+Figure 1: Examples of video-to-4D generation.
+Input: A brief video of a dynamic object, as represented by 2 frames per case; Output: Generated novel views at different timestamps.
+The generation time is also shown for each method.
+More visualized comparisons can be found in Figure 5.
 2 Related work
+[…]
 3D generation
@@
 SDS’s generality has prompted numerous subsequent efforts in both text-to-3D tasks [2, 20, 21, 22, 3, 23, 24, 25, 26, 4, 27] and image-to-3D tasks [28, 29, 30, 31] across various dimensions.
-However, SDS-based approaches often suffer from extended optimization times.
+However, SDS-based approaches often suffer from difficulty to converge and extended optimization times.
 Conversely, some efficient methods [32, 33, 7, 34, 8] have emerged.
@@
 Importantly, SyncDreamer [8] produces multi-view consistent images, offering inspiration for reconstructing 3D objects.
+[…]
 4D representation
@@
 Our work is perpendicular to all the above works, where any of them can be deployed in our reconstruction phase.
-But considering both optimization efficiency and expressive capability, we propose to represent dynamic 3D assets by a set of native 4D scene primitives.
-4D generation
-There are a few recent works dedicated for more challenging 4D object generation.
+But considering both optimization efficiency and expressive capability, we choose to represent dynamic 3D assets by a set of native 4D scene primitives, which is proved to be superior in 4DGS [15].
+[…]
+4D generation There are a few recent works dedicated for more challenging 4D object generation.
 For instance, MAV3D [13] deals with a text-to-4D problem by training Hexplane [41] with a video diffusion model and SDS loss.
@@
 To overcome this limitation, we present a novel two-staged pipeline in a generation-and-reconstruction strategy, drastically boosting the training speed by 20$\times$ whilst maintaining the quality of novel view synthesis.
-3 Method
-Figure 2: Overview of our Efficient4D approach.
-Given as the input a brief video depicting a dynamic object from a single perspectives, our model aims to generate this object with geometrical and temporal consistency under any specific view and time.
-Efficient4D comprises two components: (A) Image sequence synthesis across views and timestamps, resulting in (c) an image matrix where each row consists of multi-view geometrically consistent images and each column consists of view-specific temporally consistent images.
-(B) A novel 4D Gaussian representation model (d) that represents the scene with a number of Gaussian points.
-It can be trained efficiently and robustly under the confidence-aware (e) supervision on the generated image matrix.
-Our Efficient4D addresses the challenge of efficiently generating dynamic objects under novel views from a single-view video. As illustrated in Fig. 2, it comprises two key components:
-  • An image synthesis pipeline (Sec. 3.1) generates images across views and timestamps, ensuring sufficient geometry and temporal consistency.
-  • A novel 4D Gaussian representation model (Sec. 3.2) efficiently utilizes the synthetic images for accurate dynamic object reconstruction and novel view synthesis.
-3.1 Image synthesis across views and timestamps
-Due to the difficulty of obtaining calibrated 4D scans, our approach involves the direct generation of high-quality consistent 4D data from a single-view video which is much easier to capture (Fig.
-2(A)).
-Specifically, we seek to produce a $M\times N$ image matrix $\mathcal{D}=\{e_{ij}\}_{i,j=1}^{M,N}$ representing 2D images with geometrical and temporal consistency.
-Here, $M$ denotes timestamps, and $N$ represents views, with each matrix element corresponding to an image (Fig.
-2(c)).
-This approach combines conventional video (capturing time variation, represented by a single column in the image matrix) and 3D (capturing view variation, represented by a single row in the image matrix) generation [37, 22], offering comprehensive information for modeling a dynamic object.
-To initiate the image matrix $\mathcal{D}$, we set the first view (i.e., the first column) with $K$ frames from the input video (Fig.
-2(a)) and proceed to generate the remaining views.
-Our task involves generating multi-view consistent images from a single image for each row.
-Existing image-to-3D methods, such as SyncDreamer [22], can be leveraged for this purpose.
-However, these methods often struggle with temporal inconsistency within a specific view (i.e., continuity in the column direction) due to the independent synthesis of multi-frame images.
-To address this issue, we propose an enhanced version of SyncDreamer with improved temporal continuity, referred to as SyncDreamer-T.
-Specifically, SyncDreamer generates $N$ multi-view images $\{\mathbf{x}_{0}^{(1)},\dots,\mathbf{x}_{0}^{(N)}\}$ of a static object using a synchronized $N$-view noise predictor $\{\epsilon_{\theta}^{(n)}|n=1,\dots,N\}$ that predicts synchronized noise for noisy multi-view images $\mathbf{x}_{t}^{1:N}$.
-The noise predictor is conditioned on information correlated with all views.
-Cross-view conditioning is achieved through a spatial feature volume $\mathcal{V}\in\mathbb{R}^{F\times V\times V\times V}$ unprojected by $\mathbf{x}_{t}^{1:N}$ to inject 3D-aware features into the shared noise predictor, ensuring geometrical consistency across views for static moments.
-To impose temporal consistency, the information from $\mathcal{V}$ at different timestamps is aggregated using a time-synchronous spatial volume we design here (Fig.
-2(b)).
-A smoothing filter layer is introduced into the spatial volumes of different frames/timestamps, incorporating a weight vector $\mathbf{w}=(w_{-k},\dots,w_{0},\dots,w_{k})$.
-At each denoising step, time-synchronized spatial volumes $\tilde{\mathcal{V}}_{i}$ for each input frame $i\in\{1,2,\dots,M\}$ are constructed as:
-$\tilde{\mathcal{V}}_{i}=\sum_{j=-k}^{k}w_{i+j}\mathcal{V}_{i+j}.$ (1)
-This synchronization ensures consistent features across $k$ past and $k$ future frames during the denoising process, enhancing temporal consistency.
-With this time-synchronized spatial volumes, the proposed SyncDreamer-T is entirely training-free established on the pretrained SyncDreamer.
-To further improve temporal resolution, video frame interpolation (e.g., RIFE [12]) is applied after generating the image matrix $\mathcal{D}$ as described above.
-The midpoint interpolation is applied twice recursively, giving three additional frames between each pair of consecutive frames.
-This results in a total of $M=4K-3$ images in a column of $\mathcal{D}$.
-Refer to caption Refer to caption
-Input Feature fusion
-Figure 3: Example analysis on temporal synchronization.
-In this illustration, we manipulate the fusion ratio ($w$) across a range from 0 to 0.5 for the spatial feature volumes of two input frames.
-The results indicate that a moderate ratio can achieve superior outcomes by balancing both temporal consistency and motion independence.
-Analysis on temporal synchronization
-For a clearer insight into our temporal synchronization design, we undertake a simplified experiment involving two input frames, each with its feature volumes labeled as $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$.
-The fusion process is carried out by combining them as follows:
-$\tilde{\mathcal{V}}_{i}=(1-w)\mathcal{V}_{i}+w\mathcal{V}_{3-i},$ (2)
-where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion" as illustrated in Fig.
-3.
-For $w=0$, the original generation process is represented, where the spatial volumes of the two frames are independent, leading to temporal inconsistency.
-As $w$ approaches 0.5, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry.
-This suggests that smoothing at the feature level is effective in aligning frames over time.
-However, a challenge arises as the fused features may induce similar motion.
-While the bottom-left astronaut is stepping forward, the one at the bottom-right column does not exhibit forward motion like the top astronaut.
-Thus, a trade-off is necessary between achieving temporal consistency and preserving motion independence.
-In practical terms, it is recommended to set the ratio between 0.25 and 0.4, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry.
-Note that the results in Fig.
-3 are not sensitive to the choice in this range.
-3.2 4D generation through reconstruction
-Aiming for 3D dynamic object modeling, discrete images not suffice.
-Our next goal is to model a truly 4D content from the image matrix $\mathcal{D}$.
-For efficient modeling, we formulate a novel 4D Gaussian representation model (Fig.
-2(B)), departing from previous slow-to-train 4D reconstruction models [8].
-4D Gaussian representation model
-Our model builds upon the 3D Gaussian splatting technique introduced in [15], originally designed for static scene representation.
-We extend it to address the complexities of dynamic scenes, enabling the real-time synthesis of high-fidelity views in 4D space.
-The model represents each Gaussian defined as:
+3 Preliminary
+[…]
+4D Gaussian splatting
+The 4D Gaussian splatting (4DGS [15]) builds upon the 3D Gaussian splatting technique introduced in [16], originally designed for static scene representation.
+To address the complexities of dynamic scenes, 4DGS represents each Gaussian defined as:
 $G(\mathbf{p}|\mathbf{\mu},\Sigma)=e^{-\frac{1}{2}\left(\mathbf{p}-\mathbf{\mu}\right)^{\top}\Sigma^{-1}\left(\mathbf{p}-\mathbf{\mu}\right)},$ (1)
@@
 The 4D rotation $R$ is represented by a pair of iso-rotations, each characterized by a quaternion.
-In contrast to the 3D counterpart, our formulation incorporates the time dimension, allowing coherent modeling of dynamics in both space and time.
-Rendering
 For rendering, each Gaussian includes opacity $\alpha$ and view-dependent color $\mathbf{c}$ represented by spherical harmonics (SH).
@@
 The projection operation $P$ [16, 46] projects the world point $\mathbf{\mu}_{\mathbf{x}|t}$ to image space and transforms the covariance by ${P}(\Sigma_{\mathbf{x}|t})=JW\Sigma_{\mathbf{x}|t}W^{\top}J^{\top}$ where $W$ is the extrinsic matrix of $\mathcal{I}$ and $J$ is the Jacobian of the affine approximation of the projective transformation.
-Optimization
-In training, optimization is performed on the mean ($\mathbf{\mu}$), covariance ($\Sigma$), opacity ($\alpha$), and spherical harmonic (SH) coefficients, as well as density control including densification and pruning for each Gaussian.
-The original objective function, as presented in [15], assumes clean training data, which may not less valid for synthetic data with inherent imperfections.
-To address this, we introduce a confidence-aware objective loss formulation defined as:
-$\mathcal{L}=\lambda_{\text{rgb}}\mathcal{C}_{\text{rgb}}\mathcal{L}_{\text{rgb}}+\lambda_{\text{ssim}}\mathcal{C}_{\text{ssim}}\mathcal{L}_{\text{ssim}},$ (9)
-where $\mathcal{L}_{\text{rgb}}$ is the $L_{1}$ loss in RGB space, $\mathcal{L}_{\text{ssim}}$ is the SSIM loss [44], $\lambda_{\text{rgb/ssim}}$ is the respective weight hyper-parameter, and $\mathcal{C}_{\text{rgb/ssim}}$ is the confidence score of a generated image $\mathbf{I}$ calculated as
-$\mathcal{C}_{\text{rgb}}=1-|\mathbf{I}-\hat{\mathbf{I}}|,\quad\mathcal{C}_{\text{ssim}}=\text{SSIM}(\mathbf{I},\hat{\mathbf{I}})$ (10)
+[…]
+Score distillation sampling Score distillation sampling (SDS) is first introduced by DreamFusion [1], which is used to distill the knowledge from a pretrained diffusion model $\epsilon_{\phi}$.
+Specifically, given an image $I$ rendered from a scene representation (e.g.
+3DGS) parameterized by $\theta$, the gradient of SDS loss is calculated as:
+$\nabla_{\theta}\mathcal{L}_{\mathrm{SDS}}(\phi,I_{t})=\mathbb{E}\left[w(t)\left(\epsilon_{\phi}(I_{t};t,c)-\epsilon\right)\frac{\partial I_{t}}{\partial\theta}\right],$ (7)
+where $I_{t}$ is the perturbed image with noise $\epsilon$ at time step $t$, and $c$ is the condition (e.g. one frame of the input video in this paper).
+4 Method
+Figure 2: Overview of our Efficient4D approach.
+Given as the input (a) a brief video depicting a dynamic object from a single perspectives, our model aims to generate this object with geometrical and temporal consistency under any specific view and time.
+Efficient4D comprises two components: (A) Image sequence synthesis through (b) time-synchronous spatial volumes, resulting in (c) an image matrix where each row consists of multi-view geometrically consistent images and each column consists of view-specific temporally consistent images.
+(B) 4D Reconstruction using the generated images in (A).
+The 4D Gaussian representation can be trained efficiently and robustly under the confidence-weighted loss $\mathcal{L}_{\text{img-conf}}$ and the low-weighted SDS loss $\mathcal{L}_{SDS}$.
+Our Efficient4D addresses the challenge of efficiently generating dynamic objects under novel views from a single-view video.
+The input single-view video can be either provided by user or generated by a video generation model.
+The latter approach extends the application of our method beyond video-to-4D.
+For example, we can also achieve image-to-4D through a image-to-video diffusion model [17].
+As illustrated in Figure 2, it comprises two key components:
+  • An image synthesis pipeline (Figure 2(A)) generates images across views and timestamps, ensuring sufficient geometry and temporal consistency.
+  • An efficient and robust reconstruction process (Figure 2(B)) efficiently utilizes the synthetic images for accurate dynamic object reconstruction and novel view synthesis.
+We will elaborate on these components in Section 4.1 and 4.2, respectively.
+4.1 Image synthesis across views and timestamps
+Due to the difficulty of obtaining calibrated 4D scans, our approach involves the direct generation of high-quality consistent 4D data from a single-view video which is much easier to capture (Figure 2(a)).
+Specifically, we seek to produce a $M\times N$ image matrix $\mathcal{D}=\{e_{ij}\}_{i,j=1}^{M,N}$ representing 2D images with geometrical and temporal consistency.
+Here, $M$ denotes timestamps, and $N$ represents views, with each matrix element corresponding to an image (Figure 2(c)).
+This approach combines conventional video (capturing time variation, represented by a single column in the image matrix) and 3D (capturing view variation, represented by a single row in the image matrix) generation [47, 8], offering comprehensive information for modeling a dynamic object.
+To initiate the image matrix $\mathcal{D}$, we set the first view (i.e., the first column) with $K$ frames from the input video and proceed to generate the remaining views.
+Our task involves generating multi-view consistent images from a single image for each row.
+Existing image-to-3D methods, such as SyncDreamer [8], can be leveraged for this purpose.
+However, these methods often struggle with temporal inconsistency within a specific view (i.e., continuity in the column direction) due to the independent synthesis of multi-frame images.
+To address this issue, we propose an enhanced version of SyncDreamer with improved temporal continuity, referred to as SyncDreamer-T.
+Specifically, SyncDreamer generates $N$ multi-view images $\{\mathbf{x}_{0}^{(1)},\dots,\mathbf{x}_{0}^{(N)}\}$ of a static object using a synchronized $N$-view noise predictor $\{\epsilon_{\theta}^{(n)}|n=1,\dots,N\}$ that predicts synchronized noise for noisy multi-view images $\mathbf{x}_{t}^{1:N}$.
+The noise predictor is conditioned on information correlated with all views.
+Cross-view conditioning is achieved through a spatial feature volume $\mathcal{V}\in\mathbb{R}^{F\times V\times V\times V}$ unprojected by $\mathbf{x}_{t}^{1:N}$ to inject 3D-aware features into the shared noise predictor, ensuring geometrical consistency across views for static moments.
+To impose temporal consistency, the information from $\mathcal{V}$ at different timestamps is aggregated using a time-synchronous spatial volume we design here (Figure 2(b)).
+A smoothing filter layer is introduced into the spatial volumes of different frames/timestamps, incorporating a weight vector $\mathbf{w}=(w_{-k},\dots,w_{0},\dots,w_{k})$ which serves as the smooth filter.
+At each denoising step, time-synchronized spatial volumes $\tilde{\mathcal{V}}_{i}$ for each input frame $i\in\{1,2,\dots,M\}$ are constructed as:
+$\tilde{\mathcal{V}}_{i}=\sum_{j=-k}^{k}w_{i+j}\mathcal{V}_{i+j}.$ (8)
+This synchronization ensures consistent features across $k$ past and $k$ future frames during the denoising process, thus enhancing temporal consistency.
+With this time-synchronized spatial volumes, the proposed SyncDreamer-T is entirely training-free established on the pretrained SyncDreamer.
+To further improve temporal resolution, video frame interpolation (e.g., RIFE [48]) can be applied after generating the image matrix $\mathcal{D}$.
+The midpoint interpolation is applied twice recursively, giving three additional frames between each pair of consecutive frames.
+This results in a total of $M=4K-3$ images in a column of $\mathcal{D}$.
+Refer to caption Refer to caption
+Input Feature fusion
+Figure 3: Example analysis on temporal synchronization.
+In this illustration, we manipulate the fusion ratio ($w$) across a range from 0 to 0.5 for the spatial feature volumes of two input frames.
+The results indicate that a moderate ratio can achieve superior outcomes by balancing both temporal consistency and motion independence.
+[…]
+Analysis on temporal synchronization For a clearer insight into our temporal synchronization design, we undertake a simplified experiment involving two input frames, each with its feature volumes labeled as $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$.
+The fusion process is carried out by combining them as follows:
+$\tilde{\mathcal{V}}_{i}=(1-w)\mathcal{V}_{i}+w\mathcal{V}_{3-i},$ (9)
+where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion" as illustrated in Figure 3.
+For $w=0$, the original generation process is represented, where the spatial volumes of the two frames are independent, leading to temporal inconsistency.
+As $w$ approaches $0.5$, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry.
+This suggests that smoothing at the feature level is effective in aligning frames over time.
+However, a challenge arises as the fused features may induce similar motion.
+While the bottom-left astronaut is stepping forward, the one at the bottom-right column does not exhibit forward motion like the top astronaut.
+Thus, a trade-off is necessary between achieving temporal consistency and preserving motion independence.
+In practical terms, it is recommended to set the ratio just above $0.5$, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry.
+Also note that the results in Figure 3 are not sensitive to the choice.
+4.2 4D generation through reconstruction
+Aiming for 3D dynamic object modeling, discrete images not suffice.
+Our next goal is to model a truly 4D content from the image matrix $\mathcal{D}$.
+For efficient modeling, we have formulated the 4DGS representation model in Section 3, departing from previous slow-to-train 4D reconstruction models [40].
+[…]
+Optimization In the training of 4DGS, optimization is performed on the mean ($\mathbf{\mu}$), covariance ($\Sigma$), opacity ($\alpha$), and spherical harmonic (SH) coefficients, as well as density control including densification and pruning for each Gaussian.
+The original loss function, as presented in [15], involves both RGB loss and SSIM loss with balancing weights fixed.
+Specifically, the loss formulation is defined as:
+$\mathcal{L}_{\text{img}}=\lambda_{\text{rgb}}\mathcal{L}_{\text{rgb}}+\lambda_{\text{ssim}}\mathcal{L}_{\text{ssim}},$ (10)
+where $\mathcal{L}_{\text{rgb}}$ is the $L_{1}$ loss in RGB space, $\mathcal{L}_{\text{ssim}}$ is the SSIM loss [49], and $\lambda_{\text{rgb/ssim}}$ is the respective weight hyper-parameter.
+However, such optimization approach assumes clean training data, which may not valid for synthetic data with inherent imperfections.
+To address this, we first introduce a inconsistency-aware loss formulation with adaptive balancing weights:
+$\lambda_{\text{rgb}}=\tilde{\lambda}_{\text{rgb}}\mathcal{C}_{\text{rgb}},\quad\lambda_{\text{ssim}}=\tilde{\lambda}_{\text{ssim}}\mathcal{C}_{\text{ssim}},$ (11)
+where $\tilde{\lambda}_{\text{rgb/ssim}}$ is fixed weight and $\mathcal{C}_{\text{rgb/ssim}}$ is the adaptive confidence score of a generated image $\mathbf{I}$ calculated as
+$\mathcal{C}_{\text{rgb}}=1-|\mathbf{I}-\hat{\mathbf{I}}|,\quad\mathcal{C}_{\text{ssim}}=\text{SSIM}(\mathbf{I},\hat{\mathbf{I}}),$ (12)
 where $\hat{\mathbf{I}}$ is the unwarped image from adjacent frames estimated by optical flow.
-In such way, the confidence $\mathcal{C}_{\text{rgb/ssim}}$ function as an adaptive role in controlling the loss and gradient by assigning lower weights to inconsistent regions.
+In such way, the confidence $\mathcal{C}_{\text{rgb/ssim}}$ function as an adaptive role in controlling the loss and gradient by assigning lower weights to inconsistent regions, thus enhancing overall reconstruction quality.
+The confidence design can guarantee temporal consistency, but achieving high quality of novel view synthesis still remains challenging due to the sparsity of the generated images.
+Therefore, we also incorporate SDS loss with a relatively small weight $\lambda_{\text{SDS}}$ for smooth transition across different supervised views.
+We use the image-to-3D diffusion model [17] in SDS loss conditioned on the frame of input video at each timestamp.
+The total loss function is expressed by:
+$\mathcal{L}_{\text{total}}=\mathcal{L}_{\text{img-conf}}+\lambda_{\text{SDS}}\mathcal{L}_{\text{SDS}}$ (13)
+with $\mathcal{L}_{\text{img-conf}}=\tilde{\lambda}_{\text{rgb}}\mathcal{C}_{\text{rgb}}\mathcal{L}_{\text{rgb}}+\tilde{\lambda}_{\text{ssim}}\mathcal{C}_{\text{ssim}}\mathcal{L}_{\text{ssim}}$.
 5 Experiments
-4.1 Implementation details
-C4D [13]
-Ours
-GT
-Figure 4: Qualitative evaluation of Consistent4D (C4D) [13] and our method against ground truth (GT) on synthetic data.
-Competitors
-Video-to-4D is a relatively unexplored area, with only one existing method, Consistent4D [13].
-In comparison, we obtained the results of Consistent4D by running its released code.
-We also compare SyncDreamer [22] partially by replacing SyncDreamer-T in Sec.
-4.5.
-Evaluation videos
-To showcase the versatility of our proposed method, we conducted extensive experiments using a diverse set of data sources.
-Specifically, we focused on 36 sequences: 32 sequences released by [13] and 4 sequences processed by ourselves.
-Among the released data, seven sequences are synthetic data where ground truth are provided.
-Our four sequences are used for sparse input evaluation in Sec.
-4.4, which only contain two frames for each sequence.
+5.1 Experiment setup
+[…]
+Competitors For comparison, we mainly focus on video-to-4D and image-to-4D task.
+The competitors include Consistent4D [14] (ICLR2024), 4DGen [18] (ArXiv2024) and STAG4D [19] (ECCV2024) for video-to-4D, and Animate124 [50] (ArXiv2023) and DreamGaussian4D [51] (ArXiv2024) for image-to-4D.
+We obtained their results by running their released official code.
+We also compare SyncDreamer [8] partially by replacing SyncDreamer-T in ablation study (Section 5.6).
+[…]
+Evaluation data To showcase the versatility of our proposed method, we conducted extensive experiments using a diverse set of data sources.
+For video-to-4D, we focused on 36 sequences: 32 sequences released by [14] and 4 sequences processed by ourselves.
+Among the released data, seven sequences are synthetic data where ground truth are available.
+For image-to-4D, we used the ten images released by [50].
+Our four sequences are used for sparse input evaluation in Section 5.5, which only contain two frames for each sequence.
 Three of them, named dragon, guard, and yoxi, are rendered from 3D animated models obtained from Sketchfab [52].
 The other one named yellow face was collected from the internet.
-Each sequence consists of monocular videos captured by a stationary camera, with duration ranging from approximately 1s to 3s.
-Evaluation metrics
-As 4D generation research is still at early stage, there is no well established metric yet.
+All the data are publicly available.
+[…]
+Evaluation metrics As 4D generation research is still at early stage, there is no well established metric yet.
 However, we adopt multiple metrics by referring related works for comprehensive evaluation.
@@
 To evaluate a 4D object completely for different methods, we obtain the generated images by rendering 320 images uniformly distributed in space and time, covering 16 viewpoints and 20 timestamps.
-Implementation
-For SyncDreamer-T, we typically set the weight as $\mathbf{w}=(0.1,0.1,0.6,0.1,0.1)$ in time-synchronous spatial volume.
-The improved sampling strategy [45] can be seamlessly integrated into SyncDreamer-T.
-In the reconstruction stage, we implement a modified CUDA rasterizer tailored to 4D Gaussian.
-Gaussians are initialized randomly within the space $[-0.6,0.6]^{3}$ with identity rotations and a initial number of 20,000.
-Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 20,000 iterations with a batch size of 2.
-Gaussians are densified every 200 iterations, with densification halted at the 10,000th iteration.
-All other hyperparameters remain consistent with those in [15].
-For speed efficiency, our proposed method only costs about 2 minutes for image generation and 4 minutes for reconstruction on one A6000 GPU.
-4.2 Qualitative evaluation
-Input samples
-Refer to caption Refer to caption Refer to caption Refer to caption
-Consistent4D [13] ~165 min
-Refer to caption Refer to caption Refer to caption Refer to caption
-Our images ~2 min
-Refer to caption Refer to caption Refer to caption Refer to caption
-Our geometry ~4 min
-Refer to caption Refer to caption Refer to caption Refer to caption
-Figure 5: Qualitative comparisons. We show four images per method, captured from two viewpoints at two timestamps. Our geometry is the points at 3D mean of Gaussians.
-We first present the results on synthetic data in Fig.
-4 with ground truth shown.
-Our visual results exhibit superior accuracy in both texture and geometry when compared to the ground truth.
-For other data without ground truth, we present qualitative comparisons in Fig.
-5.
-In geometry rendering, we extract points at the mean of Gaussians.
-We also eliminate white background points and transparent points with low opacity.
-When assessing texture quality, it is important to note that both methods fall under the category of lifting 2D to 4D.
-However, Consistent4D produces watercolor-like images with low fidelity, which can be attributed to conflicts from multiple supervisory signals during prolonged optimization.
-In contrast, our method excels in directly generating high-quality 2D images.
-Subsequently, our method’s reconstruction stage yields explicit point clouds and exhibits less degradation in rendering quality compared to the SDS-based Consistent4D.
-Overall, our method consistently outperforms Consistent4D in most cases.
-4.3 Quantitative evaluation
-Table 1: Quantitative comparisons. The frame per second (FPS) is evaluated using a $256\times 256$ resolution rendering. : Measured using code released by Consistent4D [13] with one A6000 GPU.
-Consistent4D [13] Efficient4D(Ours)
-- Comparison with ground truth on synthetic data
-CLIP $\uparrow$ 0.87 0.92
-LPIPS $\downarrow$ 0.16 0.14
-- Comparison on all cases
-CLIP $\uparrow$ 0.8471 0.8721
-CLIP-T-f $\uparrow$ 0.9692 0.9702
-CLIP-T-s $\uparrow$ 0.9658 0.9726
-CLIP-T-b $\uparrow$ 0.9697 0.9707
-Generation time $\downarrow$ 120 mins 6 mins
-FPS $\uparrow$ ~20 ~1000
-Explicit geometry
-In Tab. 1, we compare Consistent4D and our method quantitatively. The metrics and data used are described in Sec. 4.1.
-We draw several observations: (1) For synthetic data, our generation results produce better CLIP and LPIPS scores.
-(2) Our method achieves superior image quality and temporal smoothness on all cases.
-(3) Our method significantly accelerates the generation process, achieving over 20$\times$ speed improvement.
-(4) Our efficient 4D Gaussian representation and splatting based rendering enable real-time rendering and provide explicit point cloud geometry.
-More specifically, the speed improvement is attributed to (1) our SDS-free design, converging optimization faster, and (2) each iteration in our method requiring much less time compared to Consistent4D.
-This time efficiency is further enhanced by avoiding recurrent evaluation in the diffusion model and prolonged forward rendering of the inefficient 4D representation as needed in the competitor at every iteration.
+[…]
+Implementation details In equation (8), the spatial volumes are smoothed locally in a sliding window style, so the consistency of generated images may be weakened as the frame number increase.
+In practice, we can adjust the weight following two rules: (i) the middle weight should be at least $0.5$; (ii) At least half of the volumes should have a positive weight.
+Here we provide some sample weights which have a good effect:
+$\mathbf{w}=\left\{\begin{matrix}(1.5,7,1.5),\quad 3\leq N<5\\ (1,1,6,1,1),\quad 5\leq N<7\\ (1,1,1,6,1,1,1),\quad 7\leq N<13\\ (1,1,1,1,1,2,14,2,1,1,1,1,1),\quad N\geq 13,\end{matrix}\right.$ (14)
+where $N$ denotes the frame number. The weight will be normalized to ensure the sum is $1$.
+In denosing process, we follow the default setting of [8] using a improved sampling strategy introduced by HarmonyView [58].
+HarmonyView redefines the score estimation by decomposing consistency and diversity.
+Please refer to [58] for details.
+In the reconstruction stage, Gaussians are initialized randomly inside the sphere with radius space 0.5 with identity rotations and a initial number of 5,000.
+Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 1500 iterations with batch size 1.
+All other 4Dgs hyperparameters remain consistent with those in [15].
+The balancing weights in loss function (equation 13) are set as $\tilde{\lambda}_{\text{rgb}}=8000,\tilde{\lambda}_{\text{ssim}}=2000,\lambda_{\text{SDS}}=1$.
+In equation (12), the calculation of estimated image $\mathbf{I}$ is implemented by frame interpolation using a optical flow based model [48].
+For each generated frame $I_{t}$, we use its adjacent four frames to interpolate the estimated frame as follows:
+$\hat{I}_{t1}=\text{Interp}(I_{t-1},I_{t+1}),~{}~{}\hat{I}_{t2}=\text{Interp}(I_{t-2},I_{t+2}).$ (15)
+Then we can compute the mean RGB confidence score:
+$\mathcal{C}_{\text{rgb}}=1-\frac{1}{2}\left(|I_{t}-\hat{I}_{t1}|+|I_{t}-\hat{I}_{t2}|\right).$ (16)
+The SSIM confidence score is similar by using equation (12).
+For speed efficiency, our proposed method only costs about 2 minutes for image generation if parallel denoising is allowed and 8 minutes for reconstruction on one A6000 GPU.
+5.2 Evaluation on synthetic data
+Figure 4: Qualitative evaluation against ground truth (GT) on synthetic data. We compare our Efficient4D with Consistent4D [14], 4DGen [18] and STAG4D [19].
+Table 1: Quantitative evaluation on synthetic data. We report CLIP and LPIPS scores between rendered images and ground truth images.
+Consistent4D [14] 4DGen [18] STAG4D [19] Efficient4D(Ours)
+CLIP $\uparrow$ 0.87 0.89 0.91 0.92
+LPIPS $\downarrow$ 0.16 0.14 0.13 0.13
+We first present the results on synthetic data in Figure 4 with ground truth shown.
+Since the data only include videos, we compare with two video-to-4D methods: Consistent4D [14], 4DGen [18] and STAG4D [19].
+Our visual results exhibit superior accuracy in both texture and geometry when compared to the ground truth, such as the direction of dinosaur’s head, the color of trump’s arm and the teeth of skull.
+The quantitative metrics are also presented in Table 1.
+The better CLIP and LPIPS scores further validate our superior generation results.
+5.3 Video-to-4D comparison
+Figure 5: Qualitative comparisons on video-to-4D generation.
+We compare our Efficient4D with Consistent4D [14], 4DGen [18] and STAG4D [19].
+For each case, we show four images per method with $0^{\circ}$ elevation.
+Our Efficient4D comprises two stages: image generation stage (Our stage-1, $30^{\circ}$ elevation) and reconstruction stage (Our stage-2, $0^{\circ}$ elevation).
+Figure 6: Qualitative comparisons on image-to-4D generation.
+We compare our Efficient4D with Animate124 [50] and DreamGaussian4D [51] (DG4D).
+For each case, we show four images per method with $0^{\circ}$ elevation.
+Our Efficient4D comprises two stages: image generation stage (Our stage-1, $30^{\circ}$ elevation) and reconstruction stage (Our stage-2, $0^{\circ}$ elevation).
+Table 2: Quantitative comparisons with state-of-the-art methods on both video-to-4D and image-to-4D generation.
+[…]
+- Video-to-4D comparison
+Consistent4D [14] 0.8471 0.9692 0.9658 0.9697 120 mins
+4DGen [18] 0.8730 0.9568 0.9568 0.9573 130 mins
+STAG4D [19] 0.8398 0.9766 0.9731 0.9760 70 mins
+Efficient4D (Ours) 0.8745 0.9766 0.9609 0.9780 10 mins
+- Image-to-4D comparison
+Animate124 [50] 0.8076 0.9673 0.9639 0.9541 540 mins
+DreamGaussian4D [51] 0.8242 0.8999 0.9072 0.9038 12 mins
+Efficient4D (Ours) 0.8350 0.9346 0.9297 0.9321 10 mins
+For data without ground truth, we present qualitative comparisons in Figure 5.
+When assessing texture quality, it is important to note that all methods fall under the category of lifting 2D to 4D.
+However, Consistent4D produce watercolor-like images with low fidelity, which can be exemplified by the blurry edges of their rendered images.
+We attribute this blur to conflicts from multiple supervisory signals during prolonged optimization.
+In contrast, our method excels in directly generating high-quality 2D images, thus providing a strong and consistent supervision in the reconstruction stage.
+Although 4DGen and STAG4D also use pseudo labels, their images are inconsistent in the temporal coordinate and SDS loss still dominates the optimization.
+Therefore, 4DGen lacks details and STAG4D suffers from floaters and blur.
+Besides, our method’s reconstruction stage supports rendering under any viewpoints while the images generated in stage-1 are discrete and sparse with a fixed elevation $30^{\circ}$.
+Overall, our method consistently outperforms the baseline methods in most cases.
+In Table 2, we compare our method with baselines quantitatively.
+The metrics and data used are described in Section 5.1.
+We draw several observations: (i) Our method achieves superior image quality and temporal smoothness on all cases validated by the higher CLIP and CLIP-T scores; (ii) Our method significantly accelerates the generation process, achieving over 10$\times$ speed improvement (10 mins vs 120 mins).
+More specifically, the speed improvement is attributed to (i) our image supervision design, converging optimization faster and reduce the training iterations, and (ii) each iteration in our method requiring much less time thanks to the efficiency of Gaussian representation.
+5.4 Image-to-4D comparison
+Next, we will evaluate our method on the image-to-4D task.
+In Figure 6, we present the visual results of three different methods.
+It can be observed that our extended method is capable of generating 4D assets with higher quality compared to those produced by state-of-the-art image-to-4D methods.
+Animate124 generates textures lacking in detail, and DreamGaussian4D tends to produce fragmented meshes.
+In contrast, the results of our Efficient4D exhibit both good geometry and high-quality textures.
+Additionally, Table 2 supports the similar conclusion that our method is superior, as seen in the video-to-4D comparison.
+Although Animate124 achieves the best CLIP-T score, This score merely indicates that the range of motion is small rather than the motion being continuous.
+From Figure 6 we can easily find that the quality of Animate124 results are lower than any other methods.
+Another notable point is that the overall CLIP scores for image-to-4D methods are lower than those for video-to-4D methods.
+This may be because the reference single-view videos in image-to-4D are generated by video diffusion models, which are inconsistent and have lower quality compared to real-world videos.
+5.5 Sparse input case
 Refer to caption Refer to caption Refer to caption
@@
 Figure 7: Given only two input frames, our method is able to generate smooth dynamics. For each case, we show three internal images from two novel views.
-4.4 Sparse input evaluation
 We assessed our method’s performance in an extremely sparse input scenario, comprising only two discrete frames.
-In such cases, we set $w=0.25$ in Eq.
-2.
-As illustrated in Fig.
-6, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency.
+In such cases, we set $w=0.25$ in equation (9).
+We also modified the code of Consistent4D [14] to make it adapt to such case.
+As illustrated in Figure 7, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency.
 In contrast, Consistent4D fails to operate effectively under such conditions.
@@
 While the toy can take different poses, it lacks autonomous movement, posing a challenge for continuous video capturing.
-In these instances, our method demonstrates effectiveness by requiring only a few key frames to produce dynamic content, thereby expanding the potential applications of the video-to-4D task.
+In these instances, our method demonstrates effectiveness by requiring only a few key frames to produce dynamic content, thereby expanding the potential applications of the 4D generation task.
 5.6 Ablation studies
+5.6.1 Ablation of image generation
 Single view
@@
 Figure 8: Ablation study on image generation, time-synchronous spatial volumes and frame interpolation. The images follow a chronological order from left to right.
-K-Plane [8] PSNR: 33.50 Time: 2 hrs
-Refer to caption Refer to caption
-No conf PSNR: 32.13 Time: 6 mins
-Refer to caption Refer to caption
-Full setting PSNR: 32.64 Time: 6 mins
-Refer to caption Refer to caption
-Reconstructed novel views
-Figure 8: Ablation study on 4D scene representation and confidence maps. The images follow a chronological order from left to right.
 Table 3: Quantitative ablation study on all our modules mesured on 4 sequences: alpaca, astronaut, rabbit and spiderman.
@@
 No interp 0.8595 0.9409 0.9443 0.9336
-No conf 0.8611 0.9658 0.9592 0.9611
 Full setting 0.8702 0.9689 0.9673 0.9684
-We performed ablation studies to assess the influence of different components in our approach.
-We compared our full method against five baseline settings: (1) Only input video are used for reconstruction; (2) Time-synchronous spatial volume is excluded; (3) Frame interpolation is excluded; (4) Confidence maps are excluded; (5) 4D Gaussian splatting is replaced with K-Plane [8].
-Note setting (2) is equivalent to the case where SyncDreamer-T is replaced with original SyncDreamer.
-The results corresponding to these settings can be found in Fig.
-7 and Fig.
-8.
+We first performed ablation studies to assess the influence of different components in our image generation stage.
+To better evaluate the image quality itself, we reconstruct the 4D Gaussian without SDS loss.
+We compared our full method against three baseline settings: (1) Only input video are used for reconstruction; (2) Time-synchronous spatial volume is excluded; (3) Frame interpolation is excluded.
+Note that setting (2) is equivalent to the case where SyncDreamer-T is replaced with original SyncDreamer.
+The results corresponding to these settings can be found in Figure 8.
 Table 3 also gives quantitative evaluations.
-Importance of synthetic data.
-We compare our generated image matrix with utilizing only input video using the same proposed 4D Gaussian representation model.
-As shown in the first row of Fig.
-7, when only relying on a single-view video, the model cannot produce any meaningful results for novel views.
+[…]
+Importance of synthetic data We compare our generated image matrix with utilizing only input video using the same 4D Gaussian representation model.
+As shown in the first row of Figure 8, when only relying on a single-view video, the model cannot produce any meaningful results for novel views.
 This indicates the importance of constructing proper training data.
-Effect of time-synchronous spatial volume.
-We assess the impact of the time-synchronous spatial volume concept introduced in SyncDreamer, as depicted in the contrast of the second and last rows in Fig.
-7 and Tab.
-2.
+[…]
+Effect of time-synchronous spatial volume We assess the impact of the time-synchronous spatial volume concept introduced in SyncDreamer, as depicted in the contrast of the second and last rows in Figure 8.
 Without time-synchronous spatial volume, the back aspects of the toy Spiderman exhibit inconsistencies, leading to distorted geometry.
-In contrast, the proposed time-synchronous spatial volume enhances both spatial and temporal consistency while preventing geometry collapse, resulting in more visually appealing image generation and higher CLIP-T scores.
-Effect of frame interpolation.
-As illustrated in the contrast between the third and last rows in Fig.
-7 and Tab.
-2, frame interpolation is effective in mitigating the blurring observed in the rendered image from novel views, thus delivering higher CLIP-T scores.
+In contrast, the proposed time-synchronous spatial volume enhances both spatial and temporal consistency while preventing geometry collapse, resulting in more visually appealing image generation and higher CLIP-T scores in Table 3.
+[…]
+Effect of frame interpolation As illustrated in the contrast between the third and last rows in Figure 8, frame interpolation is effective in mitigating the blurring observed in the rendered image from novel views, thus also delivering higher CLIP-T scores in Table 3.
 This is attributed to the low frame rates of the image matrix, which results in noticeable discontinuities.
-Effect of confidence-aware in objective.
-The integration of a confidence-aware loss in our design serves as a strategy to mitigate training data noise.
-This is evident in the contrast between the “No conf” and “Full setting” results in Fig.
-8 and Tab.
-2.
-The inclusion of confidence maps effectively reduces blurry rendering caused by inter-frame inconsistency, resulting in a significant overall improvement in quality and temporal smoothness.
-Integrating existing 4D representation model.
-Our synthetic training data is versatile and supports the optimization of 4D representation models, such as K-Plane [8].
-In the comparison shown in Fig.
-8, our 4D Gaussian splatting and K-Plane exhibit comparable proficiency in reconstructing dynamic scenes.
-Notably, our method surpasses K-Plane in speed, achieving a 20$\times$ improvement.
-Additionally, as emphasized in Sec.
-4.3, our model enables real-time rendering with explicit point cloud geometry.
+5.6.2 Ablation of reconstruction stage
+Figure 9: Ablation study on confidence maps. The images follow a chronological order from left to right.
+Figure 10: Ablation study on image supervision and SDS loss.
+In the reconstruction stage, we will study the effect of three components: (1) confidence map, (2) supervision of generated images, and (3) SDS loss. The visual results are shown in Figure 9 and 10.
+[…]
+Effect of confidence map The integration of a confidence-weighted loss in our design serves as a strategy to mitigate training data noise.
+In the left of Figure 9, we study the effect of confidence map when obvious inconsistency exists in the generated images.
+Our temporal smooth mechanism can support consistency for most areas, whilst a small fraction of lower-quality regions may introduce conflicting gradients in the reconstruction stage, thus hurting the overall quality.
+To mitigate this, our confidence aware design comes into play for weakening the supervision from those inconsistent regions.
+Since these weakly supervised regions only occupy a small proportion across all views, the missing information about texture or geometry can be compensated by the redundancy of other views and generalization ability of 4D Gaussians.
+As shown in the left of Figure 9, the inclusion of confidence maps effectively reduces blurry and inconsistent rendering even when the generated images have plausible consistency, resulting in a significant overall improvement in quality and temporal smoothness.
+[…]
+Importance of image supervision and SDS loss In Figure 10, we compare the full setting of our method with baselines without image supervision (w/o image) or SDS loss (w/o SDS).
+The absence of anchored image supervision results in bad geometry, such as the hole in the elephant ears and multiple legs of the lion.
+This can be also attributed to the conflicts of SDS loss during prolonged optimization.
+In addition, due to the sparsity of the generated images, the novel views may not be reconstructed well, leading to blur and floaters.
+By contrast, our full method avoids the disadvantages of both, being able to render clean images with good geometry.
 6 Conclusion
 This study introduces a new framework, Efficient4D, designed for generating efficiently dynamic 4D objects seamlessly from monocular videos captured by a stationary camera.
-The Efficient4D consists of two main stages: first, generating consistent multi-view videos with spatial and temporal coherence, and second, rapidly producing 4D object reconstructions using an innovative 4D Gaussian splatting model.
-Our approach, utilizing SDS-Free design and efficient 4D representation, significantly accelerates the generation process, achieving over 20 times faster speeds compared to previous work, while still delivering superior reconstruction and novel view synthesis results.
+The Efficient4D consists of two main stages: first, generating consistent multi-view videos with spatial and temporal coherence, and second, rapidly producing 4D object reconstructions.
+Our approach, utilizing image supervision with lightly weighted SDS loss, significantly accelerates the generation process, achieving about 10 times faster speeds compared to previous works, while still delivering superior reconstruction and novel view synthesis results.
 Moreover, our model is effective in extremely sparse input scenarios, requiring only two available images, thereby expanding its application scope.
-Limitations.
-Our local smoothing approach in sliding window style with our image sequence generator poses challenges in handling long-duration videos.
-To address this issue, replacing the smoothing filtering layer with a learnable attention layer featuring global receptive fields is a viable solution, particularly when dealing with extensive 4D datasets.
-Additionally, handling long videos may necessitate significant GPU memory, which can be mitigated by employing multi-GPUs or CPUs, albeit with an increase in processing time.

### v4

@@
 For example, Efficient4D takes only 10 minutes to model a dynamic object, vs 120 minutes by the previous art model Consistent4D.
-Our code is publicly available at https://github.com/fudan-zvg/Efficient4D.
+Our code and video results are publicly available at https://fudan-zvg.github.io/Efficient4D/.
 keywords:
@@
 To overcome this limitation, we present a novel two-staged pipeline in a generation-and-reconstruction strategy, drastically boosting the training speed by 20$\times$ whilst maintaining the quality of novel view synthesis.
+Recent works have attempted to generate 4D videos from a single input video.
+For example, TrajectoryCrafter [46] leverages rendered point cloud as the condition of diffusion model to facilitate novel view synthesis.
+However, the point cloud is projected from the source video and its estimated depth, so there are no points from the back view, limiting the range of novel views.
+Another work ReCamMaster [47] trains a diffusion model which introduces frame dimension conditioning for input video and camera pose conditioning for target trajectory.
+But the trajectory is still limited to the front 180.
+In contrast, our goal is to generate the complete 360 views for a dynamic object.
 3 Preliminary
@@
 The 4D rotation $R$ is represented by a pair of iso-rotations, each characterized by a quaternion.
-For rendering, each Gaussian includes opacity $\alpha$ and view-dependent color $\mathbf{c}$ represented by spherical harmonics (SH).
-Given an arbitrary view $\mathcal{I}$ defined by intrinsic and extrinsic parameters, we render the pixel at position $(u,v)$ with timestamp $t$ by blending visible Gaussians:
+[…]
 $\mathcal{I}(u,v,t)=$ $\sum_{i}f_{i}(u,v,t)\alpha_{i}\mathbf{c}_{i}(d)T_{i}$ (2)
-with $\text{with }T_{i}=$ $\prod^{i-1}_{j=1}(1-f_{j}(u,v,t)\alpha_{j}).$ (3)
+$\text{with }T_{i}=$ $\prod^{i-1}_{j=1}(1-f_{j}(u,v,t)\alpha_{j}).$ (3)
 Here, $i$ indexes the visible Gaussians sorted by depth, $d$ refers to the direction of the pixel under the view $\mathcal{I}$, and $f_{i}(u,v,t)$ denotes the influence of a Gaussian on this position.
@@
 $\tilde{\mathcal{V}}_{i}=(1-w)\mathcal{V}_{i}+w\mathcal{V}_{3-i},$ (9)
-where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion" as illustrated in Figure 3.
+where $i\in\{1,2\}$ and $w$ denotes the fusion ratio, we systematically vary the ratio $w$ from 0 to 0.5, resulting in distinct columns of “Feature fusion” as illustrated in Figure 3.
 For $w=0$, the original generation process is represented, where the spatial volumes of the two frames are independent, leading to temporal inconsistency.
-As $w$ approaches $0.5$, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry.
+As $w$ approaches 0.50.50.5, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry.
 This suggests that smoothing at the feature level is effective in aligning frames over time.
@@
 Thus, a trade-off is necessary between achieving temporal consistency and preserving motion independence.
-In practical terms, it is recommended to set the ratio just above $0.5$, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry.
+In practical terms, it is recommended to set the ratio just above 0.50.50.5, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry.
 Also note that the results in Figure 3 are not sensitive to the choice.
@@
 Implementation details In equation (8), the spatial volumes are smoothed locally in a sliding window style, so the consistency of generated images may be weakened as the frame number increase.
-In practice, we can adjust the weight following two rules: (i) the middle weight should be at least $0.5$; (ii) At least half of the volumes should have a positive weight.
+In practice, we can adjust the weight following two rules: (i) the middle weight should be at least 0.50.50.5; (ii) At least half of the volumes should have a positive weight.
 Here we provide some sample weights which have a good effect:
 $\mathbf{w}=\left\{\begin{matrix}(1.5,7,1.5),\quad 3\leq N<5\\ (1,1,6,1,1),\quad 5\leq N<7\\ (1,1,1,6,1,1,1),\quad 7\leq N<13\\ (1,1,1,1,1,2,14,2,1,1,1,1,1),\quad N\geq 13,\end{matrix}\right.$ (14)
-where $N$ denotes the frame number. The weight will be normalized to ensure the sum is $1$.
+where $N$ denotes the frame number. The weight will be normalized to ensure the sum is 111.
 In denosing process, we follow the default setting of [8] using a improved sampling strategy introduced by HarmonyView [60].
@@
 Please refer to [60] for details.
-In the reconstruction stage, Gaussians are initialized randomly inside the sphere with radius space 0.5 with identity rotations and a initial number of 5,000.
-Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 1500 iterations with batch size 1.
+In the reconstruction stage, Gaussians are initialized randomly inside the sphere with radius space 0.5 with identity rotations and a initial number of 50,000 without densification.
+Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 500 iterations with batch size 1.
 All other 4Dgs hyperparameters remain consistent with those in [15].
@@
 This may be because the reference single-view videos in image-to-4D are generated by video diffusion models, which are inconsistent and have lower quality compared to real-world videos.
-5.5 Sparse input case
+5.5 Challenging cases
 Refer to caption Refer to caption Refer to caption
@@
 Figure 7: Given only two input frames, our method is able to generate smooth dynamics. For each case, we show three internal images from two novel views.
-We assessed our method’s performance in an extremely sparse input scenario, comprising only two discrete frames.
+Figure 8: Performance on cases including rotation. We show two novel views for each case.
+We assessed our method’s performance on two challenging cases: extremely sparse input and rotation dynamics.
+In sparse case, the input comprising only two discrete frames.
 In such cases, we set $w=0.25$ in equation (9).
@@
 In these instances, our method demonstrates effectiveness by requiring only a few key frames to produce dynamic content, thereby expanding the potential applications of the 4D generation task.
+For another rotation case, we adopt the data introduced in [61].
+As shown in the first row of Figure 8, the character will rotate its body in the monocular video input, which is complex and challenging.
+However, our approach can also successfully infer the accurate geometry while Consistent4D tends to generate fragments.
+For example, from the Consistent4D results in Figure 8, we observe that the punching figure and the kicking man both suffer fragmented arms or legs.
 5.6 Ablation studies
@@
 Input/Generated images Rendered Input/Generated images Rendered
-Figure 8: Ablation study on image generation, time-synchronous spatial volumes and frame interpolation. The images follow a chronological order from left to right.
+Figure 9: Ablation study on image generation, time-synchronous spatial volumes and frame interpolation.
+For each case, we show the input images in the first row and three novel-view images generated by SyncDreamer-T at different timestamps in the other rows, followed by one rendered image from the reconstructed 4DGS at the middle timestamp.
+Each column refers to a specific timestep.
+For the row “No interp”, there is no interpolated input image in the middle, thus leading to a vacancy at timestamp $t_{2}$.
 Table 3: Quantitative ablation study on all our modules mesured on 4 sequences: alpaca, astronaut, rabbit and spiderman.
@@
 Full setting 0.8702 0.9689 0.9673 0.9684
+Figure 10: Ablation study on the temporal synchronization.
+We compare SV3D [62], SyncDreamer [8], SyncDreamer with consistent self-attention [63] (SyncDreamer-A) and our proposed Syncdreamer-T.
+We show nine consecutive frames for each case to better evalute the temporal consistency.
 We first performed ablation studies to assess the influence of different components in our image generation stage.
@@
 Table 3 also gives quantitative evaluations.
+Beyond the above components, we also ablate the design of temporal synchronization in Figure 10.
 […]
@@
 This is attributed to the low frame rates of the image matrix, which results in noticeable discontinuities.
+[…]
+Ablation on temporal synchronization We also examine the design of temporal synchronization.
+In conceptual terms, directly applying image-to-3D methods to each frame independently is inferior due to intrinsic generative randomness.
+This can be validated in Figure 10 by comparing SV3D [62], SyncDreamer [8] and SyncDremaer-T(Ours).
+The image-to-3D methods (SV3D and SyncDreamer) cannot generate a consistent shape or texture across frames.
+Therefore, a proper temporal synchronization mechanism is necessary.
+In Section 4.1, we propose SyncDreamer-T, which blends the 3D volume feature.
+We also implement a consistent self-attention mechanism introduced in [63] in SyncDreamer [8], resulting in Syncdreamer-A.
+Specifically, in each self-attention block that projects $Q_{i},K_{i},V_{i}$ corresponding to the image $I_{i}$, we randomly sample another image $I_{j}$’s $K_{j},V_{j}$, and concatenate two $K$s and $V$s as the new $K_{i}$ and $V_{i}$.
+The results in Figure 10 indicate that our method generates a more consistent texture and smoother motion than SyncDreamer-A.
+We ascribe this to the following reason: The cross-attended features may focus too much on the details that its control on global texture and motion is weakened.
 5.6.2 Ablation of reconstruction stage
@@
 Figure 12: Ablation study on image supervision and SDS loss.
-In the reconstruction stage, we will study the effect of three components: (1) confidence map, (2) supervision of generated images, and (3) SDS loss. The visual results are shown in Figure 9 and 10.
+Figure 13: Ablation study on the choice of 4D representation.
+We apply the same optimization strategy to three 4D representations: Cascade-DyNeRF [14], Deformable-3DGS [44] and 4DGS [15].
+4DGS features the fastest convergence speed.
+In the reconstruction stage, we will study the effect of three components: (1) confidence map, (2) supervision of generated images, (3) SDS loss, and (4) choice of 4D representation.
+The visual results are shown in Figure 11, 12 and 13.
 […]
@@
 By contrast, our full method avoids the disadvantages of both, being able to render clean images with good geometry.
+[…]
+Choice of 4D representation Our synthetic training data is versatile and supports the optimization of 4D representation models, such as Cascade-DyNeRF [14] and Deformable-3DGS [44] used in other 4D generation works [14, 53].
+In the comparison shown in Figure 13, 4DGS [15] converge in 500 iterations while the others do not.
+Also, thanks to the fast convergence speed of 4DGS, our method achieves significant efficiency advantage.
 6 Conclusion

### v5

@@
 To address this, we propose an efficient video-to-4D object generation framework called Efficient4D.
-It generates high-quality spacetime-consistent images under different camera views, and then uses them as labeled data to directly reconstruct the 4D content through a 4D Gaussian splatting model.
+It generates high-quality spacetime consistent images under different camera views, and then uses them as labeled data to directly reconstruct the 4D content through a 4D Gaussian splatting model.
 Importantly, our method can achieve real-time rendering under continuous camera trajectories.
@@
 Our code and video results are publicly available at https://fudan-zvg.github.io/Efficient4D/.
-keywords:
-4D Generation, Efficiency, Diffusion Model, Gaussian Splatting, Video
+journal: International Journal of Computer Vision
 1 Introduction
 Humans possess a remarkable capacity to comprehensively comprehend the spatial and temporal characteristics of a dynamic object in a brief video, even with a limited perspective, enabling them to predict its appearance in unseen viewpoints over time.
-Despite the significant advancement of 3D object generation, existing works [1, 2, 3, 4] mostly consider static scenes or objects.
-With the availability of large-scale 3D datasets [5, 6], training generalizable models capable of directly generating multi-view images becomes possible [7, 8, 9].
-These generated images can be turned into a 3D object through reconstruction techniques [10, 11].
-By further augmenting these generated static objects with predefined animations [12], dynamic 3D content can be generated.
+Despite the significant advancement of 3D object generation, existing works Poole et al.
+(2023); Lin et al.
+(2023); Chen et al.
+(2023); Wang et al.
+(2023) mostly consider static scenes or objects.
+With the availability of large-scale 3D datasets Deitke et al.
+(2023b, a), training generalizable models capable of directly generating multi-view images becomes possible Liu et al.
+(2023b, 2024); Long et al.
+(2023).
+These generated images can be turned into a 3D object through reconstruction techniques Mildenhall et al.
+(2021); Wang et al.
+(2021).
+By further augmenting these generated static objects with predefined animations 24, dynamic 3D content can be generated.
 However, this approach is still limited due to the need for fine-grained meshes as well as rigid restrictions.
-Directly generating 4D object/scene content from text description has been recently attempted [13].
-To bypass the need of exhaustively labeled training data pairs in form of (text, 4D), it trains a Neural Radiance Fields (NeRF)-like representation [10] via score distillation sampling [1] and separates the modeling of static scene and its dynamics.
+Directly generating 4D object/scene content from text description has been recently attempted Singer et al.
+(2023b).
+To bypass the need of exhaustively labeled training data pairs in form of (text, 4D), it trains a Neural Radiance Fields (NeRF)-like representation Mildenhall et al.
+(2021) via score distillation sampling Poole et al.
+(2023) and separates the modeling of static scene and its dynamics.
 Not only is this method computationally inefficient caused by heavy supervision back-propagation through a large pretrained model, but also its textual condition is highly ambiguous in expressing the intended visual content.
-In quest of the aforementioned human’s capability, a recent work [14] proposed to generate dynamic 3D object images from a single-view video (statically captured monocular video from a fixed view), namely as video-to-4D object generation.
-However, similar as [13] this method is also slow to train (e.g., 120 minutes to model a single dynamic object) in addition to complex design, hence unscalable and expensive in practice.
+In quest of the aforementioned human’s capability, a recent work Jiang et al.
+(2024) proposed to generate dynamic 3D object images from a single-view video (statically captured monocular video from a fixed view), namely as video-to-4D object generation.
+However, similar as Singer et al.
+(2023b) this method is also slow to train (e.g., 120 minutes to model a single dynamic object) in addition to complex design, hence unscalable and expensive in practice.
 To address identified limitations, we formulate an efficient video-to-4D two staged object generation method called Efficient4D.
-In the first stage, we generate spacetime-consistent images across different camera views as synthetic training data.
-This is realized by imposing temporal smoothing into a multi-view image generator (e.g., SyncDreamer [8]) in tandem with frame interpolation.
-In the second stage, we use these training data to optimize a 4D Gaussian splatting model [15].
-This is an extension of the 3D Gaussian splatting [16], originally designed for static 3D scene representation, with the temporal dimension introduced additionally, allowing for real-time rendering under continuous camera trajectories.
+In the first stage, we generate spacetime consistent images across different camera views as synthetic training data.
+This is realized by imposing temporal smoothing into a multi-view image generator (e.g., SyncDreamer Liu et al.
+(2024)) in tandem with frame interpolation.
+In the second stage, we use these training data to optimize a 4D Gaussian splatting model Yang et al.
+(2024a).
+This is an extension of the 3D Gaussian splatting Kerbl et al.
+(2023), originally designed for static 3D scene representation, with the temporal dimension introduced additionally, allowing for real-time rendering under continuous camera trajectories.
 Using Gaussian representation brings about further computational efficiency gain, when compared with NeRF based designs (Figure 1).
@@
 Along with a lightly weighted score distillation sampling loss for smooth viewpoint transitions, this technique enables robust 4D reconstruction.
-Notably, although our method takes a single-view video as input, Efficient4Dcan be easily extended to image-to-4D task by leveraging a image-to-video diffusion model [17].
-Our contributions are summarized as follows: (i) We consider for the first time the efficiency challenge with the under-studied video-to-4D object generation problem.
-(ii) We propose an efficient video-to-4D object generation pipeline, Efficient4D, characterized by directly generating high-quality training data without the need for heavy supervision back-propagation through a large pretrained model as suffered by most 3D/4D object generation approaches.
-We also extends Efficient4D to image-to-4D task.
-(iii) We introduce a inconsistency-aware confidence-weighted loss for reconstructing a 4D Gaussian splatting model using the generated training data efficiently and robustly.
-(iv) Extensive experiments on both synthetic and real data validate the significant efficiency advantage (e.g., 10$\times$ speedup) of our Efficient4D over the prior art, whilst maintaining the quality of novel view synthesis.
-Also, our method can work well under the more challenging few-shot setting where only a handful of key frames are available for training, further extending the application scope.
-Input Consistent4D [14]: 120 mins Efficient4D (Ours): 10 mins
-Input 4DGen [18]: 130 mins Efficient4D (Ours): 10 mins
-Input STAG4D [19]: 70 mins Efficient4D (Ours): 10 mins
+Notably, although our method takes a single-view video as input, Efficient4D can be easily extended to image-to-4D task by leveraging a image-to-video diffusion model Blattmann et al.
+(2023).
+[…]
+Input Consistent4D Jiang et al. (2024): 120 mins Efficient4D (Ours): 10 mins
+Input 4DGen Yin et al. (2023): 130 mins Efficient4D (Ours): 10 mins
+Input STAG4D Zeng et al. (2024): 70 mins Efficient4D (Ours): 10 mins
 Figure 1: Examples of video-to-4D generation.
@@
 2 Related work
-[…]
 3D generation
 3D generation takes two main settigns: text-to-3D and image-to-3D.
-The pioneering work, DreamFusion [1], introduces the score distillation sampling (SDS) loss for optimizing 3D shapes with diffusion models.
-SDS’s generality has prompted numerous subsequent efforts in both text-to-3D tasks [2, 20, 21, 22, 3, 23, 24, 25, 26, 4, 27] and image-to-3D tasks [28, 29, 30, 31] across various dimensions.
+The pioneering work, DreamFusion Poole et al.
+(2023), introduces the score distillation sampling (SDS) loss for optimizing 3D shapes with diffusion models.
+SDS’s generality has prompted numerous subsequent efforts in both text-to-3D tasks Lin et al.
+(2023); Metzer et al.
+(2023); Seo et al.
+(2024); Shi et al.
+(2023); Chen et al.
+(2023); Li et al.
+(2023a); Tsalicoglou et al.
+(2024); Huang et al.
+(2023); Wu et al.
+(2024b); Wang et al.
+(2023); Zhu and Zhuang (2024) and image-to-3D tasks Xu et al.
+(2023); Tang et al.
+(2023); Melas-Kyriazi et al.
+(2023); Qian et al.
+(2024) across various dimensions.
 However, SDS-based approaches often suffer from difficulty to converge and extended optimization times.
-Conversely, some efficient methods [32, 33, 7, 34, 8] have emerged.
-Notably, Point-E [32] and Shap-E [33] train models to directly generate 3D point clouds or meshes.
-Zero123 [7] focuses on generating a 2D image from an unseen view based on a single image, convertible to a 3D shape through SDS or [34].
-Importantly, SyncDreamer [8] produces multi-view consistent images, offering inspiration for reconstructing 3D objects.
-[…]
+Conversely, some efficient methods Nichol et al.
+(2022); Jun and Nichol (2023); Liu et al.
+(2023b, a, 2024) have emerged.
+Notably, Point-E Nichol et al.
+(2022) and Shap-E Jun and Nichol (2023) train models to directly generate 3D point clouds or meshes.
+Zero123 Liu et al.
+(2023b) focuses on generating a 2D image from an unseen view based on a single image, convertible to a 3D shape through SDS or Liu et al.
+(2023a).
+Importantly, SyncDreamer Liu et al.
+(2024) produces multi-view consistent images, offering inspiration for reconstructing 3D objects.
 4D representation
-Efforts to synthesize videos with free-viewpoint control in dynamic scenes have a well-documented history [35].
-For example, pre-NeRF [10] approaches challenges in reconstructing intricate scene details.
-Recent advancements in 4D representations, particularly those based on neural rendering, include D-NeRF [36], DeVRF [37] and HyperNeRF [38], which decouple geometry and motion, utilizing a canonical space and a learned deformation field.
-DynIBaR [39] deploys an image-based rendering paradigm for representing long videos with complex camera and object motions.
-Another group of methods [40, 41, 42] adopt tensor decomposition of 4D volumes to represent dynamic 3D scenes.
-Recently, Gaussian Splatting [16] has received widespread attention for its real-time high-fidelity rendering, especially its explicit point-based representation which holds great potential in modeling dynamic scenes.
+Efforts to synthesize videos with free-viewpoint control in dynamic scenes have a well-documented history Zitnick et al.
+(2004).
+For example, pre-NeRF Mildenhall et al.
+(2021) approaches challenges in reconstructing intricate scene details.
+Recent advancements in 4D representations, particularly those based on neural rendering, include D-NeRF Pumarola et al.
+(2021), DeVRF Liu et al.
+(2022) and HyperNeRF Park et al.
+(2021), which decouple geometry and motion, utilizing a canonical space and a learned deformation field.
+DynIBaR Li et al.
+(2023b) deploys an image-based rendering paradigm for representing long videos with complex camera and object motions.
+Another group of methods Fridovich-Keil et al.
+(2023); Cao and Johnson (2023); Shao et al.
+(2023) adopt tensor decomposition of 4D volumes to represent dynamic 3D scenes.
+Recently, Gaussian Splatting Kerbl et al.
+(2023) has received widespread attention for its real-time high-fidelity rendering, especially its explicit point-based representation which holds great potential in modeling dynamic scenes.
 Consequently, a significant amount of work has been proposed to explore its extension to dynamic scene modeling.
-Among them, Deformable 3D Gaussians [43] and 4DGaussian [44] integrated the deformation field with 3D Gaussian Splatting for the joint learning of the scene’s geometry and dynamics.
-SC-GS [45] represents the motion with a set of sparse control points to achieve reconstruction and motion editing.
+Among them, Deformable 3D Gaussians Yang et al.
+(2024b) and 4DGaussian Wu et al.
+(2024a) integrated the deformation field with 3D Gaussian Splatting for the joint learning of the scene’s geometry and dynamics.
+SC-GS Huang et al.
+(2024) represents the motion with a set of sparse control points to achieve reconstruction and motion editing.
 Unlike the previous radiance field-based representations often involving complex training schedules or suffering from slow convergence, these methods can be optimized efficiently, achieving real-time rendering while surpassing the past methods in terms of quality.
 Our work is perpendicular to all the above works, where any of them can be deployed in our reconstruction phase.
-But considering both optimization efficiency and expressive capability, we choose to represent dynamic 3D assets by a set of native 4D scene primitives, which is proved to be superior in 4DGS [15].
+But considering both optimization efficiency and expressive capability, we choose to represent dynamic 3D assets by a set of native 4D scene primitives, which is proved to be superior in 4DGS Yang et al.
+(2024a).
 […]
-4D generation There are a few recent works dedicated for more challenging 4D object generation.
-For instance, MAV3D [13] deals with a text-to-4D problem by training Hexplane [41] with a video diffusion model and SDS loss.
-Instead of text input, Consistent4D [14] conditions the generation of 4D object over time on a monocular video with richer and more specific information.
-However, it is computationally inefficient due to inheriting the previous SDS loss, along with complex pipeline design.
-To overcome this limitation, we present a novel two-staged pipeline in a generation-and-reconstruction strategy, drastically boosting the training speed by 20$\times$ whilst maintaining the quality of novel view synthesis.
 Recent works have attempted to generate 4D videos from a single input video.
-For example, TrajectoryCrafter [46] leverages rendered point cloud as the condition of diffusion model to facilitate novel view synthesis.
+For example, TrajectoryCrafter YU et al.
+(2025) leverages rendered point cloud as the condition of diffusion model to facilitate novel view synthesis.
 However, the point cloud is projected from the source video and its estimated depth, so there are no points from the back view, limiting the range of novel views.
-Another work ReCamMaster [47] trains a diffusion model which introduces frame dimension conditioning for input video and camera pose conditioning for target trajectory.
+Another work ReCamMaster Bai et al.
+(2025) trains a diffusion model which introduces frame dimension conditioning for input video and camera pose conditioning for target trajectory.
 But the trajectory is still limited to the front 180.
@@
 3 Preliminary
-[…]
 4D Gaussian splatting
-The 4D Gaussian splatting (4DGS [15]) builds upon the 3D Gaussian splatting technique introduced in [16], originally designed for static scene representation.
+The 4D Gaussian splatting (4DGS Yang et al.
+(2024a)) builds upon the 3D Gaussian splatting technique introduced in Kerbl et al.
+(2023), originally designed for static scene representation.
 To address the complexities of dynamic scenes, 4DGS represents each Gaussian defined as:
@@
 The 4D rotation $R$ is represented by a pair of iso-rotations, each characterized by a quaternion.
-[…]
+For rendering, each Gaussian includes opacity $\alpha$ and view-dependent color $\mathbf{c}$ represented by spherical harmonics (SH).
+Given an arbitrary view $\mathcal{I}$ defined by intrinsic and extrinsic parameters, we render the pixel at position $(u,v)$ with timestamp $t$ by blending visible Gaussians:
 $\mathcal{I}(u,v,t)=$ $\sum_{i}f_{i}(u,v,t)\alpha_{i}\mathbf{c}_{i}(d)T_{i}$ (2)
@@
 Here, $i$ indexes the visible Gaussians sorted by depth, $d$ refers to the direction of the pixel under the view $\mathcal{I}$, and $f_{i}(u,v,t)$ denotes the influence of a Gaussian on this position.
-To obtain the influence, unlike 3D Gaussian Splatting [16] which directly projects 3D Gaussians to image space, we need condition 4D Gaussians on time and then project them.
+To obtain the influence, unlike 3D Gaussian Splatting Kerbl et al.
+(2023) which directly projects 3D Gaussians to image space, we need condition 4D Gaussians on time and then project them.
 More specifically, $f(u,v,t)$ is expressed by:
@@
 $\Sigma_{\mathbf{x}|t}$ $=\Sigma_{\mathbf{x}}-\Sigma_{\mathbf{x},t}\Sigma_{t}^{-1}\Sigma_{t,\mathbf{x}}.$ (6)
-The projection operation $P$ [16, 48] projects the world point $\mathbf{\mu}_{\mathbf{x}|t}$ to image space and transforms the covariance by ${P}(\Sigma_{\mathbf{x}|t})=JW\Sigma_{\mathbf{x}|t}W^{\top}J^{\top}$ where $W$ is the extrinsic matrix of $\mathcal{I}$ and $J$ is the Jacobian of the affine approximation of the projective transformation.
-[…]
-Score distillation sampling Score distillation sampling (SDS) is first introduced by DreamFusion [1], which is used to distill the knowledge from a pretrained diffusion model $\epsilon_{\phi}$.
+The projection operation $P$ Kerbl et al. (2023); Zwicker et al. (2001) projects the world point $\mathbf{\mu}_{\mathbf{x}|t}$ to image space and transforms the covariance by
+${P}(\Sigma_{\mathbf{x}|t})=JW\Sigma_{\mathbf{x}|t}W^{\top}J^{\top}$ (7)
+where $W$ is the extrinsic matrix of $\mathcal{I}$ and $J$ is the Jacobian of the affine approximation of the projective transformation.
+Score distillation sampling Score distillation sampling (SDS) is first introduced by DreamFusion Poole et al.
+(2023), which is used to distill the knowledge from a pretrained diffusion model $\epsilon_{\phi}$.
 Specifically, given an image $I$ rendered from a scene representation (e.g.
@@
 The latter approach extends the application of our method beyond video-to-4D.
-For example, we can also achieve image-to-4D through a image-to-video diffusion model [17].
+For example, we can also achieve image-to-4D through a image-to-video diffusion model Blattmann et al.
+(2023).
 As illustrated in Figure 2, it comprises two key components:
@@
 Here, $M$ denotes timestamps, and $N$ represents views, with each matrix element corresponding to an image (Figure 2(c)).
-This approach combines conventional video (capturing time variation, represented by a single column in the image matrix) and 3D (capturing view variation, represented by a single row in the image matrix) generation [49, 8], offering comprehensive information for modeling a dynamic object.
+This approach combines conventional video (capturing time variation, represented by a single column in the image matrix) and 3D (capturing view variation, represented by a single row in the image matrix) generation Singer et al.
+(2023a); Liu et al.
+(2024), offering comprehensive information for modeling a dynamic object.
 To initiate the image matrix $\mathcal{D}$, we set the first view (i.e., the first column) with $K$ frames from the input video and proceed to generate the remaining views.
 Our task involves generating multi-view consistent images from a single image for each row.
-Existing image-to-3D methods, such as SyncDreamer [8], can be leveraged for this purpose.
+Existing image-to-3D methods, such as SyncDreamer Liu et al.
+(2024), can be leveraged for this purpose.
 However, these methods often struggle with temporal inconsistency within a specific view (i.e., continuity in the column direction) due to the independent synthesis of multi-frame images.
@@
 With this time-synchronized spatial volumes, the proposed SyncDreamer-T is entirely training-free established on the pretrained SyncDreamer.
-To further improve temporal resolution, video frame interpolation (e.g., RIFE [50]) can be applied after generating the image matrix $\mathcal{D}$.
+To further improve temporal resolution, video frame interpolation (e.g., RIFE Huang et al.
+(2022)) can be applied after generating the image matrix $\mathcal{D}$.
 The midpoint interpolation is applied twice recursively, giving three additional frames between each pair of consecutive frames.
-This results in a total of $M=4K-3$ images in a column of $\mathcal{D}$.
+This results in a total of M=4K3M=4K-3 images in a column of $\mathcal{D}$.
 Refer to caption Refer to caption
@@
 The results indicate that a moderate ratio can achieve superior outcomes by balancing both temporal consistency and motion independence.
-[…]
 Analysis on temporal synchronization For a clearer insight into our temporal synchronization design, we undertake a simplified experiment involving two input frames, each with its feature volumes labeled as $\mathcal{V}_{1}$ and $\mathcal{V}_{2}$.
@@
 For $w=0$, the original generation process is represented, where the spatial volumes of the two frames are independent, leading to temporal inconsistency.
-As $w$ approaches 0.50.50.5, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry.
+As $w$ approaches $0.5$, we observe a gradual convergence of the generated astronauts conditioned on different input images, achieving similarity in both texture and geometry.
 This suggests that smoothing at the feature level is effective in aligning frames over time.
@@
 Thus, a trade-off is necessary between achieving temporal consistency and preserving motion independence.
-In practical terms, it is recommended to set the ratio just above 0.50.50.5, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry.
+In practical terms, it is recommended to set the ratio just above $0.5$, striking a balance that ensures temporal texture consistency at a moderate cost of entangled geometry.
 Also note that the results in Figure 3 are not sensitive to the choice.
@@
 Our next goal is to model a truly 4D content from the image matrix $\mathcal{D}$.
-For efficient modeling, we have formulated the 4DGS representation model in Section 3, departing from previous slow-to-train 4D reconstruction models [40].
-[…]
+For efficient modeling, we have formulated the 4DGS representation model in Section 3, departing from previous slow-to-train 4D reconstruction models Fridovich-Keil et al.
+(2023).
 Optimization In the training of 4DGS, optimization is performed on the mean ($\mathbf{\mu}$), covariance ($\Sigma$), opacity ($\alpha$), and spherical harmonic (SH) coefficients, as well as density control including densification and pruning for each Gaussian.
-The original loss function, as presented in [15], involves both RGB loss and SSIM loss with balancing weights fixed.
+The original loss function, as presented in Yang et al.
+(2024a), involves both RGB loss and SSIM loss with balancing weights fixed.
 Specifically, the loss formulation is defined as:
 $\mathcal{L}_{\text{img}}=\lambda_{\text{rgb}}\mathcal{L}_{\text{rgb}}+\lambda_{\text{ssim}}\mathcal{L}_{\text{ssim}},$ (11)
-where $\mathcal{L}_{\text{rgb}}$ is the $L_{1}$ loss in RGB space, $\mathcal{L}_{\text{ssim}}$ is the SSIM loss [51], and $\lambda_{\text{rgb/ssim}}$ is the respective weight hyper-parameter.
+where $\mathcal{L}_{\text{rgb}}$ is the $L_{1}$ loss in RGB space, $\mathcal{L}_{\text{ssim}}$ is the SSIM loss Wang et al.
+(2004), and $\lambda_{\text{rgb/ssim}}$ is the respective weight hyper-parameter.
 However, such optimization approach assumes clean training data, which may not valid for synthetic data with inherent imperfections.
@@
 where $\tilde{\lambda}_{\text{rgb/ssim}}$ is fixed weight and $\mathcal{C}_{\text{rgb/ssim}}$ is the adaptive confidence score of a generated image $\mathbf{I}$ calculated as
-$\mathcal{C}_{\text{rgb}}=1-|\mathbf{I}-\hat{\mathbf{I}}|,\quad\mathcal{C}_{\text{ssim}}=\text{SSIM}(\mathbf{I},\hat{\mathbf{I}}),$ (12)
+[…]
 where $\hat{\mathbf{I}}$ is the unwarped image from adjacent frames estimated by optical flow.
@@
 Therefore, we also incorporate SDS loss with a relatively small weight $\lambda_{\text{SDS}}$ for smooth transition across different supervised views.
-We use the image-to-3D diffusion model [17] in SDS loss conditioned on the frame of input video at each timestamp.
+We use the image-to-3D diffusion model Blattmann et al.
+(2023) in SDS loss conditioned on the frame of input video at each timestamp.
 The total loss function is expressed by:
@@
 5.1 Experiment setup
-[…]
 Competitors For comparison, we mainly focus on video-to-4D and image-to-4D task.
-The competitors include Consistent4D [14] (ICLR2024), 4DGen [18] (ArXiv2024) and STAG4D [19] (ECCV2024) for video-to-4D, and Animate124 [52] (ArXiv2023) and DreamGaussian4D [53] (ArXiv2024) for image-to-4D.
+The competitors include Consistent4D Jiang et al.
+(2024) (ICLR2024), 4DGen Yin et al.
+(2023) (ArXiv2024) and STAG4D Zeng et al.
+(2024) (ECCV2024) for video-to-4D, and Animate124 Zhao et al.
+(2023) (ArXiv2023) and DreamGaussian4D Ren et al.
+(2023) (ArXiv2024) for image-to-4D.
 We obtained their results by running their released official code.
-We also compare SyncDreamer [8] partially by replacing SyncDreamer-T in ablation study (Section 5.6).
-[…]
+We also compare SyncDreamer Liu et al.
+(2024) partially by replacing SyncDreamer-T in ablation study (Section 5.6).
 Evaluation data To showcase the versatility of our proposed method, we conducted extensive experiments using a diverse set of data sources.
-For video-to-4D, we focused on 36 sequences: 32 sequences released by [14] and 4 sequences processed by ourselves.
+For video-to-4D, we focused on 36 sequences: 32 sequences released by Jiang et al.
+(2024) and 4 sequences processed by ourselves.
 Among the released data, seven sequences are synthetic data where ground truth are available.
-For image-to-4D, we used the ten images released by [52].
+For image-to-4D, we used the ten images released by Zhao et al.
+(2023).
 Our four sequences are used for sparse input evaluation in Section 5.5, which only contain two frames for each sequence.
-Three of them, named dragon, guard, and yoxi, are rendered from 3D animated models obtained from Sketchfab [54].
+Three of them, named dragon, guard, and yoxi, are rendered from 3D animated models obtained from Sketchfab 40.
 The other one named yellow face was collected from the internet.
 All the data are publicly available.
-[…]
 Evaluation metrics As 4D generation research is still at early stage, there is no well established metric yet.
 However, we adopt multiple metrics by referring related works for comprehensive evaluation.
-For evaluating on synthetic data, we use LPIPS score [55] and CLIP similarity [56] between rendered images and ground truth following [14].
-For the cases without ground truth, we also use CLIP-similarity between generated images and input frames as [57, 53] to measure image quality.
-For temporal smoothness, following [18, 58, 59] we use CLIP-T to measure the similarity between adjacent frames of a generated video from different views, including front (CLIP-T-f), side (CLIP-T-s) and back (CLIP-T-b) views.
+For evaluating on synthetic data, we use LPIPS score Zhang et al.
+(2018) and CLIP similarity Radford et al.
+(2021) between rendered images and ground truth following Jiang et al.
+(2024).
+For the cases without ground truth, we also use CLIP-similarity between generated images and input frames as Tang et al.
+(2024); Ren et al.
+(2023) to measure image quality.
+For temporal smoothness, following Yin et al.
+(2023); Esser et al.
+(2023); Geyer et al.
+(2023) we use CLIP-T to measure the similarity between adjacent frames of a generated video from different views, including front (CLIP-T-f), side (CLIP-T-s) and back (CLIP-T-b) views.
 To evaluate a 4D object completely for different methods, we obtain the generated images by rendering 320 images uniformly distributed in space and time, covering 16 viewpoints and 20 timestamps.
-[…]
 Implementation details In equation (9), the spatial volumes are smoothed locally in a sliding window style, so the consistency of generated images may be weakened as the frame number increase.
-In practice, we can adjust the weight following two rules: (i) the middle weight should be at least 0.50.50.5; (ii) At least half of the volumes should have a positive weight.
+In practice, we can adjust the weight following two rules: (i) the middle weight should be at least $0.5$; (ii) At least half of the volumes should have a positive weight.
 Here we provide some sample weights which have a good effect:
 $\mathbf{w}=\left\{\begin{matrix}(1.5,7,1.5),\quad 3\leq N<5\\ (1,1,6,1,1),\quad 5\leq N<7\\ (1,1,1,6,1,1,1),\quad 7\leq N<13\\ (1,1,1,1,1,2,14,2,1,1,1,1,1),\quad N\geq 13,\end{matrix}\right$. (15)
-where $N$ denotes the frame number. The weight will be normalized to ensure the sum is 111.
-In denosing process, we follow the default setting of [8] using a improved sampling strategy introduced by HarmonyView [60].
+where $N$ denotes the frame number. The weight will be normalized to ensure the sum is $1$.
+In denosing process, we follow the default setting of Liu et al.
+(2024) using a improved sampling strategy introduced by HarmonyView Woo et al.
+(2024).
 HarmonyView redefines the score estimation by decomposing consistency and diversity.
-Please refer to [60] for details.
+Please refer to Woo et al.
+(2024) for details.
 In the reconstruction stage, Gaussians are initialized randomly inside the sphere with radius space 0.5 with identity rotations and a initial number of 50,000 without densification.
 Training of 4D Gaussian Splatting is carried out using the Adam optimizer for 500 iterations with batch size 1.
-All other 4Dgs hyperparameters remain consistent with those in [15].
+All other 4DGS hyperparameters remain consistent with those in Yang et al.
+(2024a).
 The balancing weights in loss function (equation 14) are set as $\tilde{\lambda}_{\text{rgb}}=8000,\tilde{\lambda}_{\text{ssim}}=2000,\lambda_{\text{SDS}}=1$.
-In equation (12), the calculation of estimated image $\mathbf{I}$ is implemented by frame interpolation using a optical flow based model [50].
+In equation (13), the calculation of estimated image $\mathbf{I}$ is implemented by frame interpolation using a optical flow based model Huang et al.
+(2022).
 For each generated frame $I_{t}$, we use its adjacent four frames to interpolate the estimated frame as follows:
@@
 5.2 Evaluation on synthetic data
-Figure 4: Qualitative evaluation against ground truth (GT) on synthetic data. We compare our Efficient4D with Consistent4D [14], 4DGen [18] and STAG4D [19].
+Figure 4: Qualitative evaluation against ground truth (GT) on synthetic data. We compare our Efficient4D with Consistent4D Jiang et al. (2024), 4DGen Yin et al. (2023) and STAG4D Zeng et al. (2024).
 Table 1: Quantitative evaluation on synthetic data. We report CLIP and LPIPS scores between rendered images and ground truth images.
-Consistent4D [14] 4DGen [18] STAG4D [19] Efficient4D(Ours)
+Consistent4D 4DGen STAG4D Efficient4D(Ours)
 CLIP $\uparrow$ 0.87 0.89 0.91 0.92
@@
 We first present the results on synthetic data in Figure 4 with ground truth shown.
-Since the data only include videos, we compare with two video-to-4D methods: Consistent4D [14], 4DGen [18] and STAG4D [19].
+Since the data only include videos, we compare with two video-to-4D methods: Consistent4D Jiang et al.
+(2024), 4DGen Yin et al.
+(2023) and STAG4D Zeng et al.
+(2024).
 Our visual results exhibit superior accuracy in both texture and geometry when compared to the ground truth, such as the direction of dinosaur’s head, the color of trump’s arm and the teeth of skull.
@@
 Figure 5: Qualitative comparisons on video-to-4D generation.
-We compare our Efficient4D with Consistent4D [14], 4DGen [18] and STAG4D [19].
+We compare our Efficient4D with Consistent4D Jiang et al.
+(2024), 4DGen Yin et al.
+(2023) and STAG4D Zeng et al.
+(2024).
 For each case, we show four images per method with $0^{\circ}$ elevation.
@@
 Figure 6: Qualitative comparisons on image-to-4D generation.
-We compare our Efficient4D with Animate124 [52] and DreamGaussian4D [53] (DG4D).
+We compare our Efficient4D with Animate124 Zhao et al.
+(2023) and DreamGaussian4D Ren et al.
+(2023) (DG4D).
 For each case, we show four images per method with $0^{\circ}$ elevation.
@@
 - Video-to-4D comparison
-Consistent4D [14] 0.8471 0.9692 0.9658 0.9697 120 mins
-4DGen [18] 0.8730 0.9568 0.9568 0.9573 130 mins
-STAG4D [19] 0.8398 0.9766 0.9731 0.9760 70 mins
+Consistent4D Jiang et al. (2024) 0.8471 0.9692 0.9658 0.9697 120 mins
+4DGen Yin et al. (2023) 0.8730 0.9568 0.9568 0.9573 130 mins
+STAG4D Zeng et al. (2024) 0.8398 0.9766 0.9731 0.9760 70 mins
 Efficient4D (Ours) 0.8745 0.9766 0.9609 0.9780 10 mins
 - Image-to-4D comparison
-Animate124 [52] 0.8076 0.9673 0.9639 0.9541 540 mins
-DreamGaussian4D [53] 0.8242 0.8999 0.9072 0.9038 12 mins
+Animate124 Zhao et al. (2023) 0.8076 0.9673 0.9639 0.9541 540 mins
+DreamGaussian4D Ren et al. (2023) 0.8242 0.8999 0.9072 0.9038 12 mins
 Efficient4D (Ours) 0.8350 0.9346 0.9297 0.9321 10 mins
@@
 Overall, our method consistently outperforms the baseline methods in most cases.
-In Table 2, we compare our method with baselines quantitatively.
-The metrics and data used are described in Section 5.1.
-We draw several observations: (i) Our method achieves superior image quality and temporal smoothness on all cases validated by the higher CLIP and CLIP-T scores; (ii) Our method significantly accelerates the generation process, achieving over 10$\times$ speed improvement (10 mins vs 120 mins).
-More specifically, the speed improvement is attributed to (i) our image supervision design, converging optimization faster and reduce the training iterations, and (ii) each iteration in our method requiring much less time thanks to the efficiency of Gaussian representation.
+[…]
 5.4 Image-to-4D comparison
@@
 Refer to caption Refer to caption Refer to caption
-Input Ours Consistent4D [14]
+Input Ours Consistent4D
 Figure 7: Given only two input frames, our method is able to generate smooth dynamics. For each case, we show three internal images from two novel views.
@@
 In such cases, we set $w=0.25$ in equation (10).
-We also modified the code of Consistent4D [14] to make it adapt to such case.
+We also modified the code of Consistent4D Jiang et al.
+(2024) to make it adapt to such case.
 As illustrated in Figure 7, our approach successfully generates images featuring smooth motion and high spatiotemporal consistency.
@@
 In these instances, our method demonstrates effectiveness by requiring only a few key frames to produce dynamic content, thereby expanding the potential applications of the 4D generation task.
-For another rotation case, we adopt the data introduced in [61].
+For another rotation case, we adopt the data introduced in Yang et al.
+(2025).
 As shown in the first row of Figure 8, the character will rotate its body in the monocular video input, which is complex and challenging.
@@
 Refer to caption Refer to caption Refer to caption Refer to caption
+$t_{1}$ $t_{2}$ $t_{3}$
+$t_{2}$
+$t_{1}$ $t_{2}$ $t_{3}$
+$t_{2}$
 Input/Generated images Rendered Input/Generated images Rendered
@@
 Full setting 0.8702 0.9689 0.9673 0.9684
-Figure 10: Ablation study on the temporal synchronization.
-We compare SV3D [62], SyncDreamer [8], SyncDreamer with consistent self-attention [63] (SyncDreamer-A) and our proposed Syncdreamer-T.
-We show nine consecutive frames for each case to better evalute the temporal consistency.
 We first performed ablation studies to assess the influence of different components in our image generation stage.
@@
 Beyond the above components, we also ablate the design of temporal synchronization in Figure 13.
-[…]
 Importance of synthetic data We compare our generated image matrix with utilizing only input video using the same 4D Gaussian representation model.
@@
 This indicates the importance of constructing proper training data.
-[…]
 Effect of time-synchronous spatial volume We assess the impact of the time-synchronous spatial volume concept introduced in SyncDreamer, as depicted in the contrast of the second and last rows in Figure 9.
@@
 In contrast, the proposed time-synchronous spatial volume enhances both spatial and temporal consistency while preventing geometry collapse, resulting in more visually appealing image generation and higher CLIP-T scores in Table 3.
-[…]
 Effect of frame interpolation As illustrated in the contrast between the third and last rows in Figure 9, frame interpolation is effective in mitigating the blurring observed in the rendered image from novel views, thus also delivering higher CLIP-T scores in Table 3.
 This is attributed to the low frame rates of the image matrix, which results in noticeable discontinuities.
-[…]
 Ablation on temporal synchronization We also examine the design of temporal synchronization.
 In conceptual terms, directly applying image-to-3D methods to each frame independently is inferior due to intrinsic generative randomness.
-This can be validated in Figure 10 by comparing SV3D [62], SyncDreamer [8] and SyncDremaer-T(Ours).
+This can be validated in Figure 13 by comparing SV3D Voleti et al.
+(2024), SyncDreamer Liu et al.
+(2024) and SyncDremaer-T(Ours).
 The image-to-3D methods (SV3D and SyncDreamer) cannot generate a consistent shape or texture across frames.
 Therefore, a proper temporal synchronization mechanism is necessary.
-In Section 4.1, we propose SyncDreamer-T, which blends the 3D volume feature.
-We also implement a consistent self-attention mechanism introduced in [63] in SyncDreamer [8], resulting in Syncdreamer-A.
+In Section 4.1, we propose Sync-Dreamer-T, which blends the 3D volume feature.
+We also implement a consistent self-attention mechanism introduced in Zhou et al.
+(2024) in SyncDreamer Liu et al.
+(2024), resulting in Syncdreamer-A.
 Specifically, in each self-attention block that projects $Q_{i},K_{i},V_{i}$ corresponding to the image $I_{i}$, we randomly sample another image $I_{j}$’s $K_{j},V_{j}$, and concatenate two $K$s and $V$s as the new $K_{i}$ and $V_{i}$.
@@
 Figure 12: Ablation study on the choice of 4D representation.
-We apply the same optimization strategy to three 4D representations: Cascade-DyNeRF [14], Deformable-3DGS [44] and 4DGS [15].
+We apply the same optimization strategy to three 4D representations: Cascade-DyNeRF Jiang et al.
+(2024), Deformable-3DGS Wu et al.
+(2024a) and 4DGS Yang et al.
+(2024a).
 4DGS features the fastest convergence speed.
 In the reconstruction stage, we will study the effect of three components: (1) confidence map, (2) supervision of generated images, (3) SDS loss, and (4) choice of 4D representation.
-The visual results are shown in Figure 11, 12 and 13.
-[…]
-Effect of confidence map The integration of a confidence-weighted loss in our design serves as a strategy to mitigate training data noise.
+The visual results are shown in Figure 10, 11 and 12.
+Effect of confidence map The integration of a confidence weighted loss in our design serves as a strategy to mitigate training data noise.
 In the left of Figure 10, we study the effect of confidence map when obvious inconsistency exists in the generated images.
@@
 As shown in the left of Figure 10, the inclusion of confidence maps effectively reduces blurry and inconsistent rendering even when the generated images have plausible consistency, resulting in a significant overall improvement in quality and temporal smoothness.
-[…]
 Importance of image supervision and SDS loss In Figure 11, we compare the full setting of our method with baselines without image supervision (w/o image) or SDS loss (w/o SDS).
@@
 By contrast, our full method avoids the disadvantages of both, being able to render clean images with good geometry.
-[…]
-Choice of 4D representation Our synthetic training data is versatile and supports the optimization of 4D representation models, such as Cascade-DyNeRF [14] and Deformable-3DGS [44] used in other 4D generation works [14, 53].
-In the comparison shown in Figure 13, 4DGS [15] converge in 500 iterations while the others do not.
+Choice of 4D representation Our synthetic training data is versatile and supports the optimization of 4D representation models, such as Cascade-DyNeRF Jiang et al.
+(2024) and Deformable-3DGS Wu et al.
+(2024a) used in other 4D generation works Jiang et al.
+(2024); Ren et al.
+(2023).
+In the comparison shown in Figure 12, 4DGS Yang et al.
+(2024a) converge in 500 iterations while the others do not.
 Also, thanks to the fast convergence speed of 4DGS, our method achieves significant efficiency advantage.
+Figure 13: Ablation study on the temporal synchronization.
+We compare SV3D Voleti et al.
+(2024), SyncDreamer Liu et al.
+(2024), SyncDreamer with consistent self-attention Zhou et al.
+(2024) (SyncDreamer-A) and our proposed Syncdreamer-T.
+We show nine consecutive frames for each case to better evalute the temporal consistency.
 6 Conclusion
@@
 Moreover, our model is effective in extremely sparse input scenarios, requiring only two available images, thereby expanding its application scope.
+7 Data Availability Statement
+The datasets used for evaluation and comparison are available at Consistent4D Jiang et al.
+(2024) (https://github.com/yanqinJiang/Consistent4D), Animate124 Zhao et al.
+(2023) (https://github.com/HeliosZhao/Animate124/tree/main/benchmark) and Sketchfab 40 (https://sketchfab.com/).